### v2

# From Tokens to Words: On the Inner Lexicon of LLMs

Abstract
Natural language is composed of words, but modern LLMs process sub-words as input. A natural question raised by this discrepancy is whether LLMs encode words internally, and if so how. We present evidence that LLMs engage in an intrinsic detokenization process, where sub-word sequences are combined into coherent word representations. Our experiments show that this process takes place primarily within the early and middle layers of the model. They also show that it is robust to non-morphemic splits, typos and perhaps importantly—to out-of-vocabulary words: when feeding the inner representation of such words to the model as input vectors, it can “understand” them despite never seeing them during training. Our findings suggest that LLMs maintain a latent vocabulary beyond the tokenizer’s scope. These insights provide a practical, finetuning-free application for expanding the vocabulary of pre-trained models. By enabling the addition of new vocabulary words, we reduce input length and inference iterations, which reduces both space and model latency, with little to no loss in model accuracy.
1 Introduction
Large language models (LLMs) rely heavily on tokenization methods such as byte-pair encoding (BPE; Sennrich et al., 2016). Such methods often split words into multiple tokens, potentially disrupting their morphological structure (Batsuren et al., 2024). [^1] For instance, the word “unhappiness” might be tokenized into “un,” “h,” and “appiness” (see Fig. 1). Typos and other perturbations can also lead to large variations in the tokens that represent a word (Kaushal & Mahowald, 2022). Nonetheless, LLMs exhibit a remarkable ability to recover word meaning (Cao et al., 2023). This raises important questions about how models internally compose meaningful word representations from tokens, a process referred to as detokenization (Elhage et al., 2022; Gurnee et al., 2023).
In this work, we seek to understand the detokenization mechanism in LLMs. We consider two cases: words that are not part of the model BPE vocabulary, and are thus split into multiple sub-words; and single-token, in-vocabulary words that we artificially split into multiple tokens. In both cases, our experiments indicate a word-level detokenization process in LLMs, which occurs mainly in the early to middle layers. Our results hint that language models hold a latent vocabulary or inner lexicon, which they access to identify words from token sequences. [^2] This process might resemble the mental lexicon in humans (Aitchison, 2012; Marslen-Wilson et al., 1994).
We begin by examining if internal representations of token sequences reflect whether a sequence of tokens comprises a word (Section 3). We probe the model’s hidden representations (Conneau et al., 2018) of both multi-token real English words and gibberish nonwords (Frisch et al., 2000). We observe that the representations of words and nonwords substantially diverge in middle layers—a simple k-nearest neighbors classifier achieves a 89% accuracy on the task of discriminating between both groups. Overall, these results suggest that models hold a concept of recognized words.
We next explore the mechanism through which models reconstruct cohesive word representations from sub-word tokens (Section 4). To do so, we use techniques that interpret a token’s hidden states and decode them into natural language (Belrose et al., 2023; Ghandeharioun et al., 2024). We find that for both multi-token and single-token words, in most cases, the last token can be decoded as the full word after being processed by 3–5 layers, with some words requiring up to 15 layers. Interestingly, 23% of the multi-token words are never successfully decoded from their last token’s hidden states, hinting this inner lexicon does not cover all words.
Figure 1: Left: The sub-word detokenization process in LLMs. From bottom to top: (a) Tokenization and Embedding: The input string is tokenized using a sub-word tokenizer (e.g., BPE) and converted into vector embeddings; (b) Token Aggregation: the attention mechanism is used to merge sub-word tokens into the final sub-word token; (c) Word Retrieval: The model retrieves the full word representation from an internal dictionary in its feedforward (FFN) layers. This representation is added to the residual stream, until it takes over the word’s hidden representation.
Right: When taking this hidden representation and patching into another prompt, the model interprets it as the original word. In this example, the model is prompted to repeat the (single vector) hidden representation of the word “unhappiness” (originally represented as three tokens), and is able to “understand” it by regenerating the original three tokens.
We then turn to explore how models assemble these full word representations (Section 5). We first interpret the feedforward network (FFN) updates in the vocabulary space (Geva et al., 2021; Todd et al., 2024). We show that in 85% of the evaluated words, an FFN update promoting the full word’s concept is written to the last token’s residual stream (Geva et al., 2022b; Merullo et al., 2024). Importantly, this happens in the layers before the word’s representation emerges, indicating that models retrieve the reconstructed word representation from their FFN weights. Second, we compare attention patterns between standard single-token words to how multi-token words attend to their preceding sub-word tokens. We find that the last token in multi-token words attends significantly more to its previous (sub-word) tokens in layers 1–2 compared to single-token words, and significantly less in the following layers. This suggests that models initially aggregate information from previous tokens, and later almost ignore them.
Our findings have concrete applications (Section 6); they allow us to expand the model input and output BPE vocabulary with the fused representation of multi-token words found in our experiments. Importantly, this expansion requires no update to model parameters. Applying our approach to multi-token words found in English Wikipedia data (Merity et al., 2017), we find that the model successfully uses the new vocabulary entries both as inputs and during generation; its language model performance is maintained, and even slightly improves. This demonstrates the potential to dramatically reduce both input and output sequence length, and inference costs accordingly, especially in languages where the ratio of tokens per word is high (Ahia et al., 2023; Petrov et al., 2024).
Overall, our results establish that word-level detokenization is a core process in LLMs, and provide evidence of how it unfolds across model layers. Beyond improving our understanding of the internal mechanisms driving LLMs, our work lays the foundation for practical applications, particularly in terms of optimizing token management and reducing computational costs. [^3] Our code will be released here: https://github.com/schwartz-lab-NLP/Tokens2Words
2 Related Work
Tokenization
Sub-word tokenization algorithms (Wu, 2016; Kudo, 2018) are the standard for pre-processing text in modern LLMs. The most widely used method is byte pair encoding (BPE; Sennrich et al., 2016), which keeps frequent words intact and splits rare words into multiple sub-words. Recent studies propose ways to improve tokenization to ease word reconstruction (Provilkov et al., 2020; Hofmann et al., 2022; Yehezkel & Pinter, 2023; Bauwens & Delobelle, 2024) or analyze how tokenization affects model performance (Bostrom & Durrett, 2020; Church, 2020; Klein & Tsarfaty, 2020; Zouhar et al., 2023; Schmidt et al., 2024). To the best of our knowledge, we are the first to thoroughly investigate how LLMs internally reconstruct word representations.
Detokenization and stages of inference
Recent studies suggested that early LLM layers integrate local context to map raw token embeddings into coherent concepts or entities, a process called detokenization, but their observations were derived from specific case studies (Elhage et al., 2022; Lad et al., 2024). Prior work probed transformer models to explore how predictions are constructed over the course of the model, showing early layers provide local syntactic information (Tenney et al., 2019; Vulić et al., 2020; Durrani et al., 2020). Ferrando & Voita (2024) analyzed token attributions in Llama2 (Touvron et al., 2023), and observed attention heads that promote sub-word merging. Our work focuses on word-level detokenization, and goes beyond previous efforts to provide an in-depth analysis of how word representations are assembled from multiple tokens.
Interpreting the residual stream
Recent methods for interpreting the intermediate states of LLMs draw on a residual stream perspective of transformer architectures (Vaswani et al., 2017): the hidden state acts as an information stream along the layers, from which information is read at each layer, and new information is added through residual connections (Elhage et al., 2022). Thus, hidden states in any layer can be projected into the model’s vocabulary space, treating the hidden state as if it were the output of the last layer (nostalgebraist, 2020; Dar et al., 2023; Belrose et al., 2023; Yom Din et al., 2024). Similarly, Ghandeharioun et al. (2024) proposed to decode information from hidden representations into natural language, by patching (Zhang & Nanda, 2024) it into a prompt that encourages the model to verbalize the encoded information. We use both approaches to inspect how token representations evolve across layers.
LLM memories
Our hypothesis that models maintain an inner lexicon is motivated by recent work showing feedforward networks (FFN) layers in transformers act as key-value memories that encode factual and linguistic knowledge (Geva et al., 2021; Meng et al., 2022; Dai et al., 2022). Particularly, FFN layers were shown to enrich entity tokens with associated information (Meng et al., 2022; Geva et al., 2023) and promote relevant concepts in vocabulary space to build up predictions (Geva et al., 2022a), similar to our finding that word representations are pulled from FFN layers before emerging in the hidden state of the word’s last token. Concurrently to our work, Feucht et al. (2024) showed that models tend to “forget” the contents of early tokens in the last token representations of multi-token words, but not in single-token words. They also hypothesize the existence of an implicit vocabulary in LLMs, but do not further analyze the detokenization process. Combined with our results, their findings suggest that full word information retrieved from FFN layers “deletes” the existing representation rather than enriching it.
3 A Motivating Observation: LLMs Can Tell Words from Non-Words
One of our key hypotheses in this work is that LLMs hold an internal lexicon of words, which is different from the BPE lexicon. As a result, we begin by asking whether LLMs, when processing a sequence of tokens, capture some notion of whether or not this sequence forms a word.
To address this question, we construct a balanced dataset containing two groups: one with real English words, and another with artificially generated, meaningless nonwords (Frisch et al., 2000). Both groups are tokenized using the Llama2 tokenizer (Touvron et al., 2023). The word dataset consists of 10,000 distinct words sampled from the Gutenberg corpus (Gerlach & Font-Clos, 2018), with 53% of the words containing two tokens, 37.3% containing three tokens, and the rest four tokens. We generate the nonwords by shuffling tokens from the word dataset, ensuring that the prefix and suffix positions align with the original tokens’ positional probabilities. This process ensures that nonwords tokens maintain position distribution properties similar to word tokens (Fig. 2(a)). We next fit a $k$-nearest neighbors ($k$-NN) classifier ($k=4$, using Euclidean distance) on the hidden states of the last tokens of both words and nonwords, for each layer of the Llama2-7B model. The training set consists of 80% of the dataset, and the remaining 20% are used for evaluation.
(a) nonwords creation process
(b) classification results
Figure 2: Our word vs. nonword probing classification experiments: (2(a)) Dataset creation process. Top: real words from the Gutenberg corpus are tokenized using the Llama2 tokenizer. Bottom: nonwords are generated by shuffling tokens while maintaining common prefixes and suffixes; (2(b)) Classification results of words vs. nonwords: accuracy increases in early layers (3–7), peaks in the middle (13), and decreases in later layers (20–32).
Fig. 2(b) shows our probing results. Our results reveal a three-stage pattern in the model’s representation of word and nonword token sequences. In the model’s first few layers, representations from both groups are relatively indistinguishable and performance is close to chance level. Then, from layers 2 to 6, a clear distinction between the two groups emerges, until the representations remain almost completely distinct in middle layers, between layers 6 and 20. At this point, the probe achieves a stable, high accuracy, peaking at 89% on layer 13. Finally, performance slightly drops after layer 20. Importantly, word and nonword representations are much more similar in the representations of earlier tokens, as the probe only reaches 61% classification accuracy. We provide further details in Appendix A, as well as an analysis of wrongly classified words and nonwords.
Our results indicate that language models internally represent words and nonwords differently. This distinction is gradually developed in the model’s early layers and maintained throughout its middle layers. These findings support the hypothesis that the model performs a detokenization process and suggests where this process occurs. Building on these results, we next investigate how sub-word tokens are combined into word representations across model layers, and explore the internal mechanisms that facilitate this transformation.
4 Extracting Word Identity from LLM Hidden States
We have so far observed that LLMs can differentiate between words and nonwords, suggesting an internal detokenization process specific to word composition. We next dig into this process, by asking whether we can directly extract word identity from the hidden states of sub-word tokens. We start by considering single-token words (Section 4.1), and then move on to multi-token words (Section 4.2).
4.1 Single-Token Words
We first consider in-vocabulary words, which are mapped to single tokens by the tokenizer. Naively, such words don’t tell us much about detokenization, as they are represented using only one token. To address this, we artificially split them into multiple sub-word tokens. E.g., we take the single-token word “cats” and split it into two tokens: “ca” and ”ts”. We hypothesize that if the model performs detokenization, it will represent the last token of the word (“ts”) similarly to the original word token (“cats”). We iterate the WikiText-103 dataset (Merity et al., 2017) and randomly split each single-token word longer than three characters into 2–5 sub-words tokens. We then feed the model with the new sequence of tokens.
To measure the similarity between the representation of the final token and original word, we apply the logit lens method (nostalgebraist, 2020), an interpretability method that maps a hidden representation of a given token to the word whose vector in the output unembedding matrix is most closely aligned with the hidden representation. We note that this technique is typically used to inspect model predictions of the next token in intermediate layers. As we aim to identify how LLMs represent the current word, we use the input embeddings matrix instead.
We use four LLMs: Llama2-7B, Llama3-8B (Dubey et al., 2024), Mistral-7B (Jiang et al., 2023), and Yi-6B (AI et al., 2024). We report for each layer the rate of retrieval—the proportion of words for which the closest vector in the BPE space is the original (single token) word.
Our results for the Llama2-7B are shown in Fig. 3(a). [^4] Results for the other models show a similar trend, see Appendix C. The results show that starting layer at 8, the hidden state of the final token is mapped to the original word with high accuracy, which peaks at more than 80% in layer 15.555Further, we find that 93.2% of all split words are correctly mapped to the original word in at least one model layer (see Appendix D). Interestingly, afterwards this accuracy starts to decline, possibly because the model transitions into representing the prediction of the next token.
Our results indicate that LLMs perform a detokenization process in cases where a single-token word is split into multiple tokens. This process assigns the final token with a hidden representation similar to that of the full, single-token word. But how robust is this process? To address this question, we consider a different way of splitting a word into sub-tokens—by adding typos such as flipping two adjacent characters, deleting a character, and inserting a new character—see Appendix E for more details. Much like our previous experiment, this process also splits the word into multiple tokens, as corrupted words are very rarely found in BPE lexicons. We repeat the same experiment, aiming to study whether the model maps the last token to the original (correctly-spelled) word.
Our results for Llama2-7B (Fig. 3(a)) show a similar trend, though less pronounced: the model correctly matches about 40% of the corrupted words with the original, single-token, uncorrepted word. [^6] Here, 66% of the words are correctly mapped in at least one layer. We also note that the trend across layers persists, with accuracy peaking around layer 15.
(a) Rate of retrieval for single-token words
(b) Rate of retrieval for multi-tokens words
Figure 3:  (3(a)) Logit lens rate of retrieval of single-token words artificially split (blue) and split due to typos (orange). In both cases, we see an increasing rate of retrieval after the 4th layer, peaking in the middle (16–17 layers) and then dropping; (3(b)) Patchscopes rate of retrieval for multi-token words. In this case results peak in the 5th layer and then start to decline.
4.2 Multi-Token Words
We have seen that LLMs are able to detokenize single-token words that are artificially broken into sub-words into their correct (single-token) word. This indicates that these words are part of the model’s inner lexicon. But what about multi-token words? By definition, these words do not have a single embedding vector in the BPE vocabulary, so we cannot compare them to any existing vector, and thus cannot use logit lens.
Instead, we use the Patchscopes technique (Ghandeharioun et al., 2024), which interprets the hidden representation of a model using the model’s language abilities. In particular, we feed the model with the prompt “Repeat this word twice: 1) $X$ 2)”, where $X$ is the hidden representation of the last word token. Our hypothesis is that if the multi-token word is found in the model internal lexicon, than it will “understand” its representation in the input layer as well, and successfully repeat it by generating all tokens of the original word (see Fig. 1, right).
We evaluate the model by computing the rate of retrieval—the proportion of times it repeats the correct (multi-token) word. We use the same models and dataset as in the previous experiments, though in this case we do not further split or add typos, as the words are already split into multiple tokens. Particularly, we feed each multi-token word (without any context) to the model, and extract the last token’s representation from each layer. We then use Patchscopes’ prompt in a new run to check if the model regenerates the original multi-token word.
Our Llama2-7B results (Fig. 3(b)) show a striking trend: when feeding the model with vectors from layers 5–7, it is able to repeat the word in 64% of the cases, despite never seeing this vector as input before.77777.4% of the words are repeated correctly in any layer. As to the different layers, we observe a similar trend to the single-token results: retrieval rates increase rapidly at some point (here earlier than in previous experiments, around layer 3-4), reach a peak at layer 7, and then start to decline. Combined, these results suggest that the model treats multi-tokens words as if they were in the vocabulary but split into sub-word tokens, indicating a latent vocabulary that expands beyond the tokenizer’s limitations. Interestingly, we observe that 22.6% of the multi-token words are not successfully decoded from any internal layer, hinting that they might not be represented in the model’s inner lexicon.
5 How Does Detokenization Happen?
We have so far shown that LLMs perform a detokenization step: at some point, the hidden representation of the final token of a given word becomes strikingly similar to the (single-word) vector of that word. This process is robust to artificial splits of single token words, to splits due to typos, and even to multi-token words, which the model can still recognize at the input layer, despite never having seen them there during training.
We next turn to ask how does the model reconstructs full word representations from sub-word tokens? We aim to understand the dynamics of this process by analyzing the main transformer components—feedforward network layers (FFN) and the attention mechanism.
5.1 Word Retrieval using the Feedforward Mechanism
FFN layers have been shown to serve as key-value memories for storing memory (Geva et al., 2022a; 2021; Burns et al., 2024). We hypothesize that this memory might be used to store the inner lexicon as well. Particularly, we suggest that the model uses the FFNs to refine the representation of the final token, enabling it to retrieve the original word.
To test this hypothesis, we repeat the typos experiment of single-token words (Section 4.1), but this time applying logit lens to the output of the FFN layers, instead of the hidden representation. Figure 4(a) shows the retrieval rate of the original word. We observe that the retrieval of the full word concepts from the FFN layers appears a few layers earlier compared to the hidden state, which indicates that they help refine it. Nonetheless, the rate of retrieval appears to be lower in FFNs, which suggests that other factors might come into play in building the hidden representation of words. However, in Fig. 4(b) we plot the cumulative retrieval rate, i.e., for each layer, whether the word was identified in any layer so far. Our results indicate a different conclusion: the FFN update vectors match the (single-token) input representation of the word 70% of the time in layer, particularly both earlier and more frequently compared to the hidden representation. This suggest a larger variance in the exact layer in which FFNs identify the word compared to the hidden state. Our results indicate that the FFNs play a pivotal role in reconstructing word-level information from sub-word tokens.
5.2 Token Aggregation
Our results hint that LLMs retrieve words form the FFN layers, which allows the final token to represent these words. But this role of the FFN layers only starts to emerge in layers 3–4. What happens earlier? We rely on previous work that showed that in the early layers the model focuses on integrating information from nearby tokens to compose entities (Lad et al., 2024), and hypothesize that the model starts by aggregating information from the previous sub-word tokens.
(a) Word retrieval per layer
(b) Cumulative word retrieval per layer
Figure 4: Word retrieval from the FFN layers vs. from the hidden state. (4a) Retrieval rates across layers. The FFN values peak before word retrieval begins in the hidden layers. (4b) Its cumulative version, showing that word retrieval occurs earlier and more frequently in FFNs than in the hidden representation.
To test this, we take all 2-token words in a subset of 1,000 documents from WikiText-103 (a total of 5,571 words), and feed them (along with their context) to the Llama2-7B model. We then measure the average attention weights of the final token to the prefix tokens in each layer. As control, we measure the average attention weights assigned by single-token words to their preceding token.
Figure 5: Attention weights for 2-token words: Early peaks (layers 2–3) show high attention values from the 2nd sub-word token to the first, but these values decline rapidly. Attention from single-token words to their previous token shows a similar trend, though with substantially lower values at first, which become higher later.
Our results (Fig. 5) support previous findings (Lad et al., 2024)—the attention to previous tokens is high in the first 2–3 layers, but then declines sharply (up to $\sim$90decline). [^8] Experiments with 3- and 4-token words show a similar trend, see Appendix B. Considering the attention from single-word tokens to their previous token, we also observe this peak phenomenon, but importantly—here the initial peak is far lower than in our multi-token words [^9] The diverging patterns between single and multi-token words are statistically significant; see Appendix B.. However, we also note that the following values are in turn higher. These results suggest that LLM strongly attend to their sub-word tokens at first, but then largely ignore them.
Our results suggest that the model performs a detokenization process by first aggregating information from the prefix tokens into the final token’s hidden representation. It then refines the representation of the final token in the FFN layers to retrieve the full word. This two-stage process of token aggregation and concept retrieval provides insight into the internal mechanisms of LLMs and how they handle sub-word tokens to reconstruct word-level representations.
6 Expanding LLM Vocabulary without Finetuning
We have shown that language models internally fuse multi-token words into single-token representations, which models can decode to generate the complete word. This raises the question whether models can make more general use of such fused representations when processing textual input, by using them in place of the original, multi-token embeddings. Conversely, models were shown to foreshadow several future tokens in a single hidden state without being explicitly trained to do so (Pal et al., 2023); can we leverage the fused representations to enable models to generate multi-token words in a single inference step?
We next explore whether such vocabulary expansion is possible without any updates to the model’s parameters. [^10] We do add new entries to the model’s embedding and unembedding matrices, but these are inherently required to expand the vocabulary and represent the new vocabulary words. The potential of such methods is high, as even tokenizers intentionally trained for multilingual support result in extremely longer tokenization lengths for non-English languages (up to 13x for some languages), affecting cost and latency (Ahia et al., 2023; Sengupta et al., 2023; Petrov et al., 2024). Still, previous attempts to expand tokenizer vocabulary are limited in number, and require further training on large corpora (Kim et al., 2024; Zhao et al., 2024) or attempt to force alignment between languages in embedding space (Wang et al., 2019), rather than rely on the internal detokenization mechanism of LLMs.
We first experiment with expanding the input embeddings matrix $E$, leaving the output unembedding matrix $U$ unchanged. To extract fused representations, we begin with the same approach as in Section 4.2: Given a multi-token word, we run it through the model using a prompt template $P$ [^11] We use the template “x x x x” where x is the new vocabulary word; preliminary experiments showed it provides better results than the template “x”, possibly due to the effect a token’s position has on its representation. We then identify the first layer in which the last token is successfully decoded into the full word (Ghandeharioun et al., 2024). We extract the hidden state of the last token in that layer, and add it as the word’s fused representation to $E$. If the word is never successfully decoded from the representation in any of the layers, we don’t attempt to add it to the vocabulary, as we assume it is not in the model’s inner lexicon. We experiment with two approaches: identity, where we use the fused representation as is; and linear, where we first transform it with a linear map $T_{\ell,E}$. We learn $T_{\ell,E}$ by fitting a linear regression from the layer $\ell$ hidden states of all single-token words when passed with the template $P$, to their corresponding input embeddings. Importantly, $T_{\ell,E}$ is both not part of the original model, and does not depend on the new, originally multi-token words we add to the vocabulary.
Original vocabulary tokens Newly added words
Top-1 Acc. Mean Reciprocal Rank Top-1 Acc. Mean Reciprocal Rank
Original vocabulary 57.7% 0.679 47.5% 0.556
Last token embedding ($E$) 54.5% 0.647
Fused-Identity ($E$) 53.5% 0.638
Fused-Linear ($E$) 56.7% 0.669
Fused-Linear ($E$ + $U$) 56.7% 0.667 19.5% 0.598
Table 1: Performance of finetuning-free vocabulary expansion for Llama2-7B on WikiText-103. Different rows correspond to baselines (first two rows) and proposed approaches (last three). Methods with $E$ add new multi-token words as single-token entries in the embedding matrix, while those with $U$ add new entries to the unembedding matrix. Column groups correspond to evaluations of language model performance on different token groups: those already in the vocabulary (left) and the newly added entries (right, relevant only when expanding $U$). Results marked with $\ast$ are for predicting the 1st token of a multi-token word (rather than predicting the new entry of the complete word). Our best approach (Fused-Linear) leverages the fused representations emerging within models after detokenization, to add new vocabulary entries for multi-token words. We show it maintains performance on the original vocabulary, and enables the model to successfully utilize entries in $U$ that were never observed during training.
We apply our approach to expand the vocabulary of Llama2-7B with all multi-token words found in the WikiText-103 test set, resulting in 4.3k newly added words. To examine whether the model can correctly use the new entries, we evaluate it in a language modeling setup, when all newly added words are replaced with their fused single-token embeddings, similar to our setup in Section 4.2 (see righthand side in Fig. 1). We compare our approach to two baselines: using the original vocabulary without any new entries; and using the unprocessed last token embedding of each new word as its new vocabulary entry (instead of the fused representations). For evaluation, we cannot use perplexity, the standard LM measure, as the different approaches have different vocabularies, which influences perplexity scores and prevents a fair comparison. Instead, we measure the token-level top-1 accuracy and the mean reciprocal rank of the ground truth tokens; naturally, we do not report the ranking of sub-word tokens forming the newly added words, as they only serve as ground truth in our baseline.
We report our results in Table 1. Examining model performance when using the fused representations in the identity approach, we find that while the model can successfully decode the full word information when explicitly prompted to do so, it fails to utilize the fused representation for inference, and achieves slightly worse performance than when using the unprocessed last token embeddings. However, using the linear approach by first applying $T_{\ell,E}$ transforms the fused representation into a usable form for initial layers, as we find the model can then utilize the newly added words and maintains its performance. [^12] A possible explanation for the importance of $T_{\ell,E}$ is that input embeddings have very different distributions from intermediate hidden states, as LLMs are extremely sensitive to 1st layer ablations (Lad et al., 2024).
Given the model’s ability to use fused representations as inputs, we next enable the model to generate multi-token words in a single inference step, by adding new entries to the unembedding matrix $U$ derived from their fused representations. In light of our previous results, we only experiment with the linear approach. We follow the same steps as before, this time applying linear maps $T_{\ell,U}$ which we learn by fitting linear regressions from the layer $\ell$ hidden states of all existing single-token words, to their corresponding rows in the unembedding matrix. We evaluate whether the model chooses to generate the new vocabulary entries, or rather the first token of the original multi-token words (as it does when using the unedited vocabulary).
Surprisingly, we find the model assigns, on average, higher rank to the new vocabulary entries over their original 1st sub-word tokens in the appropriate contexts (Table 1, last row). While the top-1 accuracy is higher for the 1st sub-word tokens in the unedited vocabulary, we find that when new vocabulary entries are not assigned high ranks, the original sub-word will often be predicted. We further observe that in 7.1% of examples, the new entry achieves the top-1 rank even though the original sub-word in the baseline does not. Notably, this also does not compromise performance on existing vocabulary words. Our results suggest that the model can sometimes try to generate the fused representation, but as it does not exist in its (original) unembedding vocabulary, the first sub-word token is selected instead. Overall, our findings show potential to substantially decrease inference times for LLMs. In future work, we will explore their application to multilingual LLMs (Petrov et al., 2024).
7 Conclusion
The ability of LLMs to comprehend and generate language relies on intricate internal processes, and understanding these mechanisms is crucial for improving model performance as well as making them more efficient. In this work, we unraveled the word detokenization process, shedding light on how models internally transform fragmented sub-word tokens into coherent word representations formed at the last token. Our results indicate that this mechanism manifests in early to middle layers, where models attempt to reconstruct words by mapping them to an inner lexicon using their FFN layers. We provided evidence this lexicon is more exhaustive than the tokenizer’s vocabulary, and could facilitate the model’s ability to recognize words even amidst noise.
Our work also unlocks practical avenues for optimizing tokenization, as well as the speed and cost of inference. We demonstrated one such application and presented a finetuning-free method to expand the vocabulary of LLMs. We hope our work paves the way for more efficient and versatile models.
Acknowledgements
This work was supported in part by the Israel Science Foundation (grant no. 2045/21) and by NSF-BSF grant 2020793. In addition, we thank Iddo Yosha for his valuable feedback.
Appendix A Words vs. Nonwords Analysis
To further analyze results from Section 3, we conduct a failure analysis, focusing on false negatives (FN) and false positives (FP). The former, valid words misclassified as gibberish, are often rare or complex words, suggesting that the model’s internal vocabulary may lack representations for less common words. On the other hand, false positives—where gibberish is misclassified as a word—typically involved nonwords that closely resemble valid words, likely due to shared sub-word structures. See Table 2 for a few examples.
(a) accuracy postfix and last token
(b) positive predictions over prefix and last token
Figure 6: Comparisons between last token and the prefixes. (Left) we see that the accuracy over the prefix tokens reaching only 61% vs the 89% with the last token’s representation. (Right) a more detailed statistics over the positive predictions. The TP rate over the last token’s representations reaching 84% while reaching only 43% for the prefix ones
Original Word Predicted Value True Value Status
_Unitarianism gibberish word FN
_killy gibberish word FN
_quadruropic word gibberish FP
_nonwith word gibberish FP
Table 2: Examples of false negative (FN) and false positive (FP) for the word vs. nonword experiment (Section 3).
Appendix B Statistical Analysis of Token Aggregation in Multi-token Words
We repeat the multi-token experiments Section 5.2 for 3- and 4-token words. Results Fig. 7 show a very similar trend to the 2-token words (Fig. 5).
We present the detailed statistical analysis conducted of our experiments in Section 5.2 to examine token aggregation in multi-token words compared to single-token words (control group). The objective is to investigate whether the model exhibits significantly different attention patterns between these groups, indicative of the detokenization process.
Layer p-value (Multi $>$ Single) p-value (Single $>$ Multi) Significance
0 $9.999\times 10^{-1}$ $5.891\times 10^{-5}$ Single-token $>$ Multi-token (***)
1 $0.000$ $1.000$ Multi-token $>$ Single-token (***)
2 $0.000$ $1.000$ Multi-token $>$ Single-token (***)
3 $1.000$ $1.252\times 10^{-273}$ Single-token $>$ Multi-token (***)
4 $1.000$ $0.000$ Single-token $>$ Multi-token (***)
5 $1.000$ $4.466\times 10^{-242}$ Single-token $>$ Multi-token (***)
6 $1.000$ $0.000$ Single-token $>$ Multi-token (***)
7 $1.000$ $0.000$ Single-token $>$ Multi-token (***)
8 $1.000$ $0.000$ Single-token $>$ Multi-token (***)
9 $1.000$ $0.000$ Single-token $>$ Multi-token (***)
10 $1.000$ $0.000$ Single-token $>$ Multi-token (***)
11 $1.000$ $7.847\times 10^{-87}$ Single-token $>$ Multi-token (***)
12 $1.000$ $6.307\times 10^{-303}$ Single-token $>$ Multi-token (***)
13 $1.000$ $1.417\times 10^{-255}$ Single-token $>$ Multi-token (***)
14 $1.000$ $2.397\times 10^{-172}$ Single-token $>$ Multi-token (***)
15 $1.000$ $0.000$ Single-token $>$ Multi-token (***)
16 $1.000$ $5.857\times 10^{-25}$ Single-token $>$ Multi-token (***)
17 $1.000$ $1.414\times 10^{-152}$ Single-token $>$ Multi-token (***)
18 $5.646\times 10^{-5}$ $9.999\times 10^{-1}$ Multi-token $>$ Single-token (***)
19 $1.000$ $7.919\times 10^{-161}$ Single-token $>$ Multi-token (***)
20 $1.000$ $2.036\times 10^{-286}$ Single-token $>$ Multi-token (***)
21 $1.399\times 10^{-190}$ $1.000$ Multi-token $>$ Single-token (***)
22 $1.000$ $3.647\times 10^{-44}$ Single-token $>$ Multi-token (***)
23 $1.000$ $6.008\times 10^{-13}$ Single-token $>$ Multi-token (***)
24 $1.000$ $1.062\times 10^{-15}$ Single-token $>$ Multi-token (***)
25 $5.274\times 10^{-68}$ $1.000$ Multi-token $>$ Single-token (***)
26 $1.000$ $3.670\times 10^{-12}$ Single-token $>$ Multi-token (***)
27 $0.000$ $1.000$ Multi-token $>$ Single-token (***)
28 $1.000$ $2.446\times 10^{-10}$ Single-token $>$ Multi-token (***)
29 $1.438\times 10^{-202}$ $1.000$ Multi-token $>$ Single-token (***)
30 $9.506\times 10^{-99}$ $1.000$ Multi-token $>$ Single-token (***)
31 $6.731\times 10^{-1}$ $3.270\times 10^{-1}$ Not significant (ns)
Table 3: Results of one-sided t-tests comparing attention weights between multi-token and single-token words across layers over Llama2-7B model.
To determine significance, we performed two one-sided t-tests per layer for the Llama2-7B case: (1) Testing whether attention to prefix tokens in multi-token words is higher than to previous tokens in single-token words. (2) Testing whether attention to previous tokens in single-token words is higher than to prefix tokens in multi-token words.
Table 3 shows the p-values and significance levels for each layer. Significance levels are denoted as ns (not significant), and *** ($p<0.001$).
(a) Llama2-7B Attention weights for 3-tokens words
(b) Llama2-7B Attention weights for 4-tokens words
Figure 7: Analysis for 3- and 4-token words for Llama2-7B. The higher attention pattern in layers 1-2, while lower values are observed afterwards, is consistent with our results for 2-token words (Fig. 5).
The results reveal that in layers 1 and 2, attention to prefix tokens in multi-token words is significantly higher than in single-token words, suggesting the early phase of Token Aggregation. From Layers 3 to 17, attention to single-token words is higher, indicating a shift in attention focus from prefix attention in relation regular close tokens. Notably, there are intermittent increases in attention to prefix tokens at Layers 18, 21, 25, 27, 29, and 30, possibly signaling a transfer into a prediction ensembling phase in which the attention in general is less important (Artzy & Schwartz, 2024) and therefore we don’t see a coherent pattern.
In conclusion, the attention mechanisms differ significantly between multi-token and single-token words, confirming that the detokenization process involves initial amplification of attention to sub-word tokens followed by a reduction as the model processes word representations.
Appendix C Word Retrieval for Single-Token and Multi-Token Words Across All Models Comparison
This section presents a detailed comparison of word retrieval performance (Section 4) across several models (Llama2-7B, Llama3-8B, Yi-6B, and Mistral-7B) for both single-token and multi-token words. The evaluation focuses on how effectively each model retrieves the original word across layers, especially in challenging cases like artificially separated single-token words, typos, and multi-token words.
Figure 8: Layer-wise word retrieval rates for single-token and multi-token words across all models.
Across all experiments, we observe a similar trend where the retrieval performance increases over the first several layers, peaks around the middle layers, and then decreases in the later layers. The main difference across models lies in the peak performance, especially in cases involving typos and multi-token words, where more advanced models such as Llama3-8B and Mistral-7B demonstrate superior performance in reconstructing the original word representations.
Appendix D Cumulative Word Retrieval for Single-Token and Multi-Token Words
In this section, we present the results of cumulative word retrieval (Section 4.1) across various models, focusing on both single-token and multi-token words. For each model, we analyze the ability of the LLM to retrieve the original word from sub-word tokens across its layers. The cumulative retrieval is calculated as the proportion of words that are successfully retrieved at each layer, with the percentage increasing as more words are recovered throughout the model’s layers.
Single-Token Words: The cumulative word retrieval for single-token words—those that are artificially split into sub-word tokens (via typographical errors or manual splits)—shows a rapid increase in retrieval success in the early layers. For Llama2-7B, for instance, cumulative retrieval reaches 93.2% for words split by manual intervention, and 66% for words affected by typos by the middle layers. This pattern is observed across models, with retrieval generally peaking around layers 15-20.
Multi-Token Words: For multi-token words, which are naturally split due to being out-of-vocabulary for the tokenizer, the cumulative retrieval process follows a similar trajectory. However, in models like Llama2-7B, the retrieval peaks earlier in the model, with a cumulative retrieval rate of 77.41%. Other models like Llama3 and Yi show higher cumulative retrieval rates, suggesting improved efficiency in handling multi-token words, potentially due to larger model capacities and internal dictionaries.
The similarity in cumulative retrieval between single-token and multi-token words suggests that LLMs treat out-of-vocabulary words in a manner similar to sub-word-tokenized words, accessing a latent vocabulary to reconstruct full word representations.
Figure 9: Cumulative word retrieval for single-token and multi-token words across all models.
Appendix E Introducing Typos for Single-Token Words
In this section, we describe the process of introducing typos into single-token words to split them into multiple tokens (Section 4.1). The modification applies to words longer than four characters and involves randomly performing one of three operations: substituting two characters, deleting a character, or inserting a new character. By introducing these slight variations, the word becomes unfamiliar to the tokenizer, causing it to be divided into multiple smaller tokens during tokenization. Particularly, this process results in splitting words into 2–5 tokens. Table 4 shows examples of the different splits.
Description perturbed . new tokens
Substitution of two characters develpoment [‘de’, ‘vel’, ‘p’, ‘oment’]
Deletion of one character develoment [‘de’, ‘vel’, ‘oment’]
Insertion of one character develfopment [‘dev’, ‘elf’, ‘op’, ‘ment’]
Table 4: Examples of the different typos we consider, exemplified by perturbing the single-token word “development” (Section 4.1).

### v3

@@
-From Tokens to Words: On the Inner Lexicon of LLMs
+From Tokens to Words: On Detokenization and the Inner Lexicon of LLMs
 Abstract
-Natural language is composed of words, but modern LLMs process sub-words as input.
+Natural language is composed of words, but modern large language models (LLMs) process sub-words as input.
 A natural question raised by this discrepancy is whether LLMs encode words internally, and if so how.
-We present evidence that LLMs engage in an intrinsic detokenization process, where sub-word sequences are combined into coherent word representations.
-Our experiments show that this process takes place primarily within the early and middle layers of the model.
-They also show that it is robust to non-morphemic splits, typos and perhaps importantly—to out-of-vocabulary words: when feeding the inner representation of such words to the model as input vectors, it can “understand” them despite never seeing them during training.
+We present evidence that LLMs engage in an intrinsic detokenization process, where sub-word sequences are combined into coherent whole-word representations at their last token.
+Our experiments show that this process primarily takes place within the early and middle layers of the model.
+We further demonstrate its robustness to arbitrary splits (e.g., “cats” to “ca” and “ts”), typos, and importantly—to out-of-vocabulary words: when feeding the last token internal representations of such words to the model as input, it can “understand” them as the complete word despite never seeing such representations as input during training.
 Our findings suggest that LLMs maintain a latent vocabulary beyond the tokenizer’s scope.
@@
 By enabling the addition of new vocabulary words, we reduce input length and inference iterations, which reduces both space and model latency, with little to no loss in model accuracy.
+[^1] We release our code at https://github.com/schwartz-lab-NLP/Tokens2Words.
 1 Introduction
@@
 [^2] For instance, the word “unhappiness” might be tokenized into “un,” “h,” and “appiness” (see Fig.
-1).
-Typos and other perturbations can also lead to large variations in the tokens that represent a word (Kaushal & Mahowald, 2022).
+1) Typos and other perturbations can also lead to large variations in the tokens that represent a word (Kaushal & Mahowald, 2022).
 Nonetheless, LLMs exhibit a remarkable ability to recover word meaning (Cao et al., 2023).
@@
 In this work, we seek to understand the detokenization mechanism in LLMs.
-We consider two cases: words that are not part of the model BPE vocabulary, and are thus split into multiple sub-words; and single-token, in-vocabulary words that we artificially split into multiple tokens.
+We consider two cases: words that are not part of the model’s BPE vocabulary, and are thus split into multiple sub-words; and single-token, in-vocabulary words that we artificially split into multiple tokens.
 In both cases, our experiments indicate a word-level detokenization process in LLMs, which occurs mainly in the early to middle layers.
@@
 [^3] This process might resemble the mental lexicon in humans (Aitchison, 2012; Marslen-Wilson et al., 1994).
-We begin by examining if internal representations of token sequences reflect whether a sequence of tokens comprises a word (Section 3).
+We begin by examining if internal representations of token sequences reflect whether or not a sequence of tokens comprises a word (Sec.
+3).
 We probe the model’s hidden representations (Conneau et al., 2018) of both multi-token real English words and gibberish nonwords (Frisch et al., 2000).
-We observe that the representations of words and nonwords substantially diverge in middle layers—a simple k-nearest neighbors classifier achieves a 89% accuracy on the task of discriminating between both groups.
+We observe that the representations of words and nonwords substantially diverge in middle layers—a simple k-nearest neighbors classifier achieves a 89% accuracy on the task of discriminating between the two groups.
 Overall, these results suggest that models hold a concept of recognized words.
-We next explore the mechanism through which models reconstruct cohesive word representations from sub-word tokens (Section 4).
+We next explore the mechanism through which models reconstruct cohesive word representations from sub-word tokens (Sec.
+4).
 To do so, we use techniques that interpret a token’s hidden states and decode them into natural language (Belrose et al., 2023; Ghandeharioun et al., 2024).
@@
 Figure 1: Left: The sub-word detokenization process in LLMs.
-From bottom to top: (a) Tokenization and Embedding: The input string is tokenized using a sub-word tokenizer (e.g., BPE) and converted into vector embeddings; (b) Token Aggregation: the attention mechanism is used to merge sub-word tokens into the final sub-word token; (c) Word Retrieval: The model retrieves the full word representation from an internal dictionary in its feedforward (FFN) layers.
+From bottom to top: (a) Tokenization and Embedding: The input string is tokenized using a sub-word tokenizer (e.g., BPE) and converted into vector embeddings; (b) Token Aggregation: The attention mechanism relays information from the word’s previous tokens (“un”, “h”) into its final sub-word representation (“appiness”); (c) Word Retrieval: The model retrieves the full word representation from an implicit internal lexicon in its feedforward (FFN) layers.
 This representation is added to the residual stream, until it takes over the word’s hidden representation.
-Right: When taking this hidden representation and patching into another prompt, the model interprets it as the original word.
+Right: When taking this hidden representation and patching it into another prompt, the model interprets it as the original word.
 In this example, the model is prompted to repeat the (single vector) hidden representation of the word “unhappiness” (originally represented as three tokens), and is able to “understand” it by regenerating the original three tokens.
-We then turn to explore how models assemble these full word representations (Section 5).
+We then turn to explore how models assemble these full word representations (Sec.
+5).
 We first interpret the feedforward network (FFN) updates in the vocabulary space (Geva et al., 2021; Todd et al., 2024).
-We show that in 85% of the evaluated words, an FFN update promoting the full word’s concept is written to the last token’s residual stream (Geva et al., 2022b; Merullo et al., 2024).
-Importantly, this happens in the layers before the word’s representation emerges, indicating that models retrieve the reconstructed word representation from their FFN weights.
+We show that in 85% of the evaluated words, an FFN update promoting the full word’s concept is written to the last token’s residual stream (Geva et al., 2022; Merullo et al., 2024).
+Importantly, this happens in the layers before the full word’s representation emerges, indicating that models retrieve the reconstructed word representation from their FFN weights.
 Second, we compare attention patterns between standard single-token words to how multi-token words attend to their preceding sub-word tokens.
-We find that the last token in multi-token words attends significantly more to its previous (sub-word) tokens in layers 1–2 compared to single-token words, and significantly less in the following layers.
+We find that the last token in multi-token words attends significantly more to its previous tokens (which are sub-words of the same word) in layers 1–2 compared to single-token words (where these previous tokens represent other words), and significantly less in the following layers.
 This suggests that models initially aggregate information from previous tokens, and later almost ignore them.
-Our findings have concrete applications (Section 6); they allow us to expand the model input and output BPE vocabulary with the fused representation of multi-token words found in our experiments.
+See Fig.
+1 for an overview.
+Our findings have concrete applications (Sec.
+6); they allow us to expand the model input and output BPE vocabulary with the fused representation of multi-token words found in our experiments.
 Importantly, this expansion requires no update to model parameters.
-Applying our approach to multi-token words found in English Wikipedia data (Merity et al., 2017), we find that the model successfully uses the new vocabulary entries both as inputs and during generation; its language model performance is maintained, and even slightly improves.
-This demonstrates the potential to dramatically reduce both input and output sequence length, and inference costs accordingly, especially in languages where the ratio of tokens per word is high (Ahia et al., 2023; Petrov et al., 2024).
+Applying our approach to multi-token words found in English Wikipedia data (Merity et al., 2017), we find that the model successfully uses the new vocabulary entries both as inputs and during generation: its language model performance is maintained, and even slightly improves.
+This demonstrates the potential to dramatically reduce both input and output sequence length, and inference costs accordingly, especially in languages where the ratio of tokens per word is high (Ahia et al., 2023; Petrov et al., 2023).
 Overall, our results establish that word-level detokenization is a core process in LLMs, and provide evidence of how it unfolds across model layers.
-Beyond improving our understanding of the internal mechanisms driving LLMs, our work lays the foundation for practical applications, particularly in terms of optimizing token management and reducing computational costs.
-[^3] Our code will be released here: https://github.com/schwartz-lab-NLP/Tokens2Words
+Beyond improving our understanding of the internal mechanisms driving LLMs, our work lays the foundation for practical applications, particularly in optimizing token management and reducing computational costs.
 2 Related Work
@@
 Sub-word tokenization algorithms (Wu, 2016; Kudo, 2018) are the standard for pre-processing text in modern LLMs.
-The most widely used method is byte pair encoding (BPE; Sennrich et al., 2016), which keeps frequent words intact and splits rare words into multiple sub-words.
-Recent studies propose ways to improve tokenization to ease word reconstruction (Provilkov et al., 2020; Hofmann et al., 2022; Yehezkel & Pinter, 2023; Bauwens & Delobelle, 2024) or analyze how tokenization affects model performance (Bostrom & Durrett, 2020; Church, 2020; Klein & Tsarfaty, 2020; Zouhar et al., 2023; Schmidt et al., 2024).
+The most widely used method is byte-pair encoding (BPE; Sennrich et al., 2016), which keeps frequent words intact and splits rare ones into multiple sub-words.
+Recent studies proposed ways to improve tokenization to consider word structure (Provilkov et al., 2020; Hofmann et al., 2022; Yehezkel & Pinter, 2023; Bauwens & Delobelle, 2024) or analyze how tokenization affects model performance (Bostrom & Durrett, 2020; Church, 2020; Klein & Tsarfaty, 2020; Zouhar et al., 2023; Schmidt et al., 2024).
 To the best of our knowledge, we are the first to thoroughly investigate how LLMs internally reconstruct word representations.
 Detokenization and stages of inference
-Recent studies suggested that early LLM layers integrate local context to map raw token embeddings into coherent concepts or entities, a process called detokenization, but their observations were derived from specific case studies (Elhage et al., 2022; Lad et al., 2024).
-Prior work probed transformer models to explore how predictions are constructed over the course of the model, showing early layers provide local syntactic information (Tenney et al., 2019; Vulić et al., 2020; Durrani et al., 2020).
-Ferrando & Voita (2024) analyzed token attributions in Llama2 (Touvron et al., 2023), and observed attention heads that promote sub-word merging.
+Early LLM layers have been shown to integrate local context and map raw token embeddings into representations of concepts or entities—a process called detokenization.
+However, such observations were based on specific case studies (Elhage et al., 2022; Lad et al., 2024).
+More generally, recent work showed early layers provide local syntactic information (Tenney et al., 2019; Vulić et al., 2020; Durrani et al., 2020; Sajjad et al., 2022) or focus on extracting information from previous tokens (Ben Artzy & Schwartz, 2024).
+Ferrando & Voita (2024) analyzed token attributions in LLMs, observing attention heads that promote sub-word merging (Correia et al., 2019).
 Our work focuses on word-level detokenization, and goes beyond previous efforts to provide an in-depth analysis of how word representations are assembled from multiple tokens.
 Interpreting the residual stream
-Recent methods for interpreting the intermediate states of LLMs draw on a residual stream perspective of transformer architectures (Vaswani et al., 2017): the hidden state acts as an information stream along the layers, from which information is read at each layer, and new information is added through residual connections (Elhage et al., 2022).
-Thus, hidden states in any layer can be projected into the model’s vocabulary space, treating the hidden state as if it were the output of the last layer (nostalgebraist, 2020; Dar et al., 2023; Belrose et al., 2023; Yom Din et al., 2024).
+Recent methods for interpreting the intermediate states of LLMs draw on a residual stream perspective: the hidden state acts as an information stream along the layers, from which information is read at each layer, and new information is added through residual connections (Elhage et al., 2022).
+Thus, hidden states at any layer can be projected into the model’s vocabulary space, treating the hidden state as if it were the output of the last layer (nostalgebraist, 2020; Dar et al., 2023; Belrose et al., 2023; Yom Din et al., 2024).
 Similarly, Ghandeharioun et al.
@@
 LLM memories
-Our hypothesis that models maintain an inner lexicon is motivated by recent work showing feedforward networks (FFN) layers in transformers act as key-value memories that encode factual and linguistic knowledge (Geva et al., 2021; Meng et al., 2022; Dai et al., 2022).
-Particularly, FFN layers were shown to enrich entity tokens with associated information (Meng et al., 2022; Geva et al., 2023) and promote relevant concepts in vocabulary space to build up predictions (Geva et al., 2022a), similar to our finding that word representations are pulled from FFN layers before emerging in the hidden state of the word’s last token.
-Concurrently to our work, Feucht et al.
-(2024) showed that models tend to “forget” the contents of early tokens in the last token representations of multi-token words, but not in single-token words.
-They also hypothesize the existence of an implicit vocabulary in LLMs, but do not further analyze the detokenization process.
-Combined with our results, their findings suggest that full word information retrieved from FFN layers “deletes” the existing representation rather than enriching it.
+The idea of an inner lexicon aligns with recent work showing feedforward networks (FFN) layers in transformers act as key-value memories that encode factual and linguistic knowledge (Geva et al., 2021; Meng et al., 2022b; Dai et al., 2022).
+Particularly, FFNs were shown to enrich entity tokens with associated information (Meng et al., 2022b; Geva et al., 2023) and promote relevant concepts in vocabulary space to build up predictions (Geva et al., 2022).
+Our work expands on these findings, showing that word representations are pulled from FFN layers before emerging in the hidden state of the word’s last token.
+Inner lexicon structure
+Inspired by studies on how concepts are encoded in LLMs (Park et al., 2024; 2025) and gradually promoted throughout their layers (Geva et al., 2022; Merullo et al., 2024), we consider the inner lexicon a “soft” lexicon, which (a) combines multiple vectors to form word representations (rather than a key-value dictionary); and (b) is not unique, i.e., a word might be stored and retrieved in more than one layer.
+Concurrently, Feucht et al.
+(2024) present evidence for an implicit vocabulary in LLMs, showing that models “forget” preceding tokens in multi-token words or multi-word entities, but remember previous tokens when processing single-token words.
 3 A Motivating Observation: LLMs Can Tell Words From Non-Words
 One of our key hypotheses in this work is that LLMs hold an internal lexicon of words, which is different from the BPE lexicon.
-As a result, we begin by asking whether LLMs, when processing a sequence of tokens, capture some notion of whether or not this sequence forms a word.
+We begin by asking whether LLMs, when processing a sequence of tokens, capture some notion of whether or not this sequence forms a word.
 To address this question, we construct a balanced dataset containing two groups: one with real English words, and another with artificially generated, meaningless nonwords (Frisch et al., 2000).
@@
 We generate the nonwords by shuffling tokens from the word dataset, ensuring that the prefix and suffix positions align with the original tokens’ positional probabilities.
-This process ensures that nonwords tokens maintain position distribution properties similar to word tokens (Fig.
+For example, the token “ing”, extracted from the final position of real words in the dataset (e.g., “directing” tokenized as “direct”, “ing”), is retained as a suffix in the nonwords dataset.
+This process ensures that nonword tokens maintain position distribution properties similar to word tokens, preserving natural positional patterns and mitigating potential distributional biases (Fig.
 2(a)).
-We next fit a $k$-nearest neighbors ($k$-NN) classifier ($k=4$, using Euclidean distance) on the hidden states of the last tokens of both words and nonwords, for each layer of the Llama2-7B model.
+We next apply a $k$-nearest neighbors ($k$-NN) probing classifier ($k=4$, using Euclidean distance) on the hidden states of the last tokens of both words and nonwords, for each layer of the Llama2-7B model.
 The training set consists of 80% of the dataset, and the remaining 20% are used for evaluation.
-(a) nonwords creation process
+(a) Nonword creation process
 (b) Classification results
 Figure 2: Our word vs.
-nonword probing classification experiments: (2(a)) Dataset creation process.
-Top: real words from the Gutenberg corpus are tokenized using the Llama2 tokenizer.
-Bottom: nonwords are generated by shuffling tokens while maintaining common prefixes and suffixes; (2(b)) Classification results of words vs.
-nonwords: accuracy increases in early layers (3–7), peaks in the middle (13), and decreases in later layers (20–32).
-Fig.
-2(b) shows our probing results.
-Our results reveal a three-stage pattern in the model’s representation of word and nonword token sequences.
-In the model’s first few layers, representations from both groups are relatively indistinguishable and performance is close to chance level.
-Then, from layers 2 to 6, a clear distinction between the two groups emerges, until the representations remain almost completely distinct in middle layers, between layers 6 and 20.
+nonword probing classification experiments.
+(2(a)) Dataset creation process.
+Top: words from the Gutenberg corpus are tokenized using the Llama2 tokenizer.
+Bottom: nonwords are generated by shuffling tokens while maintaining their positions within the word; (2(b)Classification results of words vs.
+nonwords.
+Using the last token shows strong results (reaching up to 90% accuracy), which sharply rise in early layers (3–7), peak in the middle (13), and decrease in later layers (20–32).
+Using the penultimate token shows substantially lower scores, suggesting that the high classification accuracy is specifically tied to the presence of a complete word rather than token co-occurrence patterns.
+Our results (Fig.
+2(b), blue) reveal a three-stage pattern in the model’s representation of word and nonword token sequences.
+In the model’s first few layers, representations from both groups are relatively indistinguishable and accuracy is close to chance level.
+Then, from layers 2 to 6, a clear distinction between the two groups emerges, until the representations are almost completely separate in middle layers, between layers 6 and 20.
 At this point, the probe achieves a stable, high accuracy, peaking at 89% on layer 13.
-Finally, performance slightly drops after layer 20.
-Importantly, word and nonword representations are much more similar in the representations of earlier tokens, as the probe only reaches 61% classification accuracy.
-We provide further details in Appendix A, as well as an analysis of wrongly classified words and nonwords.
-Our results indicate that language models internally represent words and nonwords differently.
+Finally, accuracy slightly drops after layer 20.444We perform a similar experiment using a dataset of morphologically plausible nonwords (ARC Nonword Database; Rastle et al.
+2002), which shows a similar trend.
+See App.
+A.
+Our results indicate the LLMs can distinguish between words and nonwords.
+But this distinction can be attributed to the distributional properties of words: it might be the case that models are recognizing common sequences of tokens, rather than identifying whole words.
+To test this hypothesis, we repeat the same experiment, this time using the penultimate token representation, for words three tokens or longer.
+By definition, such words also as frequently co-occur with the initial sub-word tokens in their words as the final tokens.
+Our results (Fig.
+2(b), orange line) show that the probe only reaches 61% classification accuracy, indicating that the high classification accuracy does not stem from token co-occurrence, but is tied to the presence of a whole word.
+See App.
+A for an analysis of misclassified words and nonwords.
+Overall, our results show that language models internally represent words and nonwords differently.
 This distinction is gradually developed in the model’s early layers and maintained throughout its middle layers.
@@
 We next dig into this process, by asking whether we can directly extract word identity from the hidden states of sub-word tokens.
-We start by considering single-token words (Section 4.1), and then move on to multi-token words (Section 4.2).
+We start by considering single-token words (Sec.
+4.1), and then move on to multi-token words (Sec.
+4.2).
 4.1 Single-Token Words
@@
 To address this, we artificially split them into multiple sub-word tokens.
-E.g., we take the single-token word “cats” and split it into two tokens: “ca” and ”ts”.
+For example, we take the single-token word “cats” and split it into two tokens: “ca” and ”ts”.
 We hypothesize that if the model performs detokenization, it will represent the last token of the word (“ts”) similarly to the original word token (“cats”).
 We iterate the WikiText-103 dataset (Merity et al., 2017) and randomly split each single-token word longer than three characters into 2–5 sub-words tokens.
-We then feed the model with the new sequence of tokens.
+We then feed the model the new sequence of tokens preceded by the last 100 tokens that came before each split word in the original text as context.
+[^5] We observe similar trends, in this and in other experiments in this section, when passing the split words without their context, albeit with slightly lower rates of retrieval.
 To measure the similarity between the representation of the final token and original word, we apply the logit lens method (nostalgebraist, 2020), an interpretability method that maps a hidden representation of a given token to the word whose vector in the output unembedding matrix is most closely aligned with the hidden representation.
 We note that this technique is typically used to inspect model predictions of the next token in intermediate layers.
-As we aim to identify how LLMs represent the current word, we use the input embeddings matrix instead.
-We use four LLMs: Llama2-7B, Llama3-8B (Dubey et al., 2024), Mistral-7B (Jiang et al., 2023), and Yi-6B (AI et al., 2024).
-We report for each layer the rate of retrieval—the proportion of words for which the closest vector in the BPE space is the original (single token) word.
+As we aim to identify how LLMs represent the current word, we use the input embeddings matrix rather than the unembedding one.
+[^6] We also compute a second measure, cosine similarity, to compare between the hidden representations and input embeddings.
+We find similar results to those obtained using logit lens.
+For details, see App.
+F.
+We study four LLMs: Llama2-7B, Llama3-8B (Dubey et al., 2024), Mistral-7B (Jiang et al., 2023), and Yi-6B (AI et al., 2024).
+For each layer, we report the rate of retrieval—the proportion of words for which the closest vector in the vocabulary space is the original (single token) word.
 Our results for the Llama2-7B are shown in Fig.
 3(a).
-[^4] Results for the other models show a similar trend, see Appendix C.
-The results show that starting layer at 8, the hidden state of the final token is mapped to the original word with high accuracy, which peaks at more than 80% in layer 15.555Further, we find that 93.2% of all split words are correctly mapped to the original word in at least one model layer (see Appendix D).
-Interestingly, afterwards this accuracy starts to decline, possibly because the model transitions into representing the prediction of the next token.
+[^7] Results for the other models (App.
+C) show a similar trend.
+Starting at layer 8, the hidden state of the final token is mapped to the original word with high accuracy, which peaks at more than 80% in layer 15.
+Interestingly, this accuracy then starts to decline, possibly because the model transitions into representing the prediction of the next token.
+We also find that 93.2% of all split words are correctly mapped to the original word in at least one model layer (see Sec.
+C.1).
 Our results indicate that LLMs perform a detokenization process in cases where a single-token word is split into multiple tokens.
@@
 But how robust is this process?
-To address this question, we consider a different way of splitting a word into sub-tokens—by adding typos such as flipping two adjacent characters, deleting a character, and inserting a new character—see Appendix E for more details.
-Much like our previous experiment, this process also splits the word into multiple tokens, as corrupted words are very rarely found in BPE lexicons.
-We repeat the same experiment, aiming to study whether the model maps the last token to the original (correctly-spelled) word.
+To address this question, we consider a different way of splitting a word into sub-tokens—adding typos.
+For each word, we randomly flip two adjacent characters, delete a character, or insert a new character (see App.
+D for more details).
+Much like our previous experiment, this process splits the word into multiple tokens, as corrupted words are rarely found in BPE lexicons.
+We repeat the same experiment as above, aiming to study whether models map the last token to the original (correctly-spelled) word.
 Our results for Llama2-7B (Fig.
-3(a)) show a similar trend, though less pronounced: the model correctly matches about 40% of the corrupted words with the original, single-token, uncorrepted word.
+3(a)) show a similar trend, though less pronounced: the model correctly matches about 40% of corrupted words with the original, single-token, uncorrupted word.
 [^8] Here, 66% of the words are correctly mapped in at least one layer.
@@
 (b) Rate of retrieval for multi-tokens words
-Figure 3:  (3(a)) Logit lens rate of retrieval of single-token words artificially split (blue) and split due to typos (orange).
+Figure 3:  (3(a)) Logit lens rate of retrieval of single-token words artificially split (blue line) and split due to typos (orange).
 In both cases, we see an increasing rate of retrieval after the 4th layer, peaking in the middle (16–17 layers) and then dropping; (3(b)) Patchscopes rate of retrieval for multi-token words.
@@
 4.2 Multi-Token Words
-We have seen that LLMs are able to detokenize single-token words that are artificially broken into sub-words into their correct (single-token) word.
+We have seen that LLMs are able to detokenize single-token words that are artificially broken into sub-words back into their correct (single-token) word.
 This indicates that these words are part of the model’s inner lexicon.
@@
 In particular, we feed the model with the prompt “Repeat this word twice: 1) $X$ 2)”, where $X$ is the hidden representation of the last word token.
-Our hypothesis is that if the multi-token word is found in the model internal lexicon, than it will “understand” its representation in the input layer as well, and successfully repeat it by generating all tokens of the original word (see Fig.
+Our hypothesis is that if the multi-token word is found in the model’s internal lexicon, then it will “understand” its representation in the input layer as well, and successfully repeat it by generating all tokens of the original word (see Fig.
 1, right).
-We evaluate the model by computing the rate of retrieval—the proportion of times it repeats the correct (multi-token) word.
+We evaluate the model by computing the rate of retrieval—the proportion of times it generates the correct (multi-token) word.
 We use the same models and dataset as in the previous experiments, though in this case we do not further split or add typos, as the words are already split into multiple tokens.
-Particularly, we feed each multi-token word (without any context) to the model, and extract the last token’s representation from each layer.
+Particularly, we feed each multi-token word (along with its 100-tokens context, as before) to the model, and extract the last token’s representation from each layer.
 We then use Patchscopes’ prompt in a new run to check if the model regenerates the original multi-token word.
 Our Llama2-7B results (Fig.
-3(b)) show a striking trend: when feeding the model with vectors from layers 5–7, it is able to repeat the word in 64% of the cases, despite never seeing this vector as input before.77777.4% of the words are repeated correctly in any layer.
+3(b)) show a striking trend: when feeding the model with vectors from layers 5–7, it is able to repeat the word in 64% of the cases, despite never seeing this vector as input before.
+[^9] Overall, 77.4% of the words are repeated correctly in at least one layer.
 As to the different layers, we observe a similar trend to the single-token results: retrieval rates increase rapidly at some point (here earlier than in previous experiments, around layer 3-4), reach a peak at layer 7, and then start to decline.
 Combined, these results suggest that the model treats multi-tokens words as if they were in the vocabulary but split into sub-word tokens, indicating a latent vocabulary that expands beyond the tokenizer’s limitations.
-Interestingly, we observe that 22.6% of the multi-token words are not successfully decoded from any internal layer, hinting that they might not be represented in the model’s inner lexicon.
+Interestingly, we observe that 22.6% of the multi-token words are never successfully decoded from any internal layer, hinting that they might not be represented in the model’s inner lexicon.
 5 How Does Detokenization Happen?
-We have so far shown that LLMs perform a detokenization step: at some point, the hidden representation of the final token of a given word becomes strikingly similar to the (single-word) vector of that word.
+We have so far shown that LLMs perform a detokenization step: at some point, the hidden representation of the final token of a given word becomes strikingly similar to the (single-tokenx) vector of that word.
 This process is robust to artificial splits of single token words, to splits due to typos, and even to multi-token words, which the model can still recognize at the input layer, despite never having seen them there during training.
-We next turn to ask how does the model reconstructs full word representations from sub-word tokens?
-We aim to understand the dynamics of this process by analyzing the main transformer components—feedforward network layers (FFN) and the attention mechanism.
+We next turn to ask how does the model reconstruct full word representations from sub-word tokens?
+We aim to understand the dynamics of this process by analyzing the main transformer components: feedforward network layers (FFN) and the attention mechanism.
 5.1 Word Retrieval Using the Feedforward Mechanism
-FFN layers have been shown to serve as key-value memories for storing memory (Geva et al., 2022a; 2021; Burns et al., 2024).
+FFN layers have been shown to serve as key-value tables for storing memory (Geva et al., 2022; 2021; Meng et al., 2022b).
 We hypothesize that this memory might be used to store the inner lexicon as well.
 Particularly, we suggest that the model uses the FFNs to refine the representation of the final token, enabling it to retrieve the original word.
-To test this hypothesis, we repeat the typos experiment of single-token words (Section 4.1), but this time applying logit lens to the output of the FFN layers, instead of the hidden representation.
+(a) Word retrieval per layer
+(b) Cumulative word retrieval per layer
+Figure 4: Word retrieval in FFN layers vs.
+in the hidden states.
+(4a) Retrieval rates across layers.
+The FFN values peak before word retrieval begins in the hidden layers.
+(4b) Its cumulative version, showing word retrieval occurs earlier and more frequently in FFNs than in the hidden representation.
+To test this hypothesis, we repeat the typos experiment of single-token words (Sec.
+4.1), but this time applying logit lens to the output of the FFN layers, instead of the hidden representation.
 Figure 4(a) shows the retrieval rate of the original word.
-We observe that the retrieval of the full word concepts from the FFN layers appears a few layers earlier compared to the hidden state, which indicates that they help refine it.
+We observe that the retrieval of the full word concepts from the FFN layers occurs a few layers earlier compared to the hidden state, which indicates that they help refine it.
 Nonetheless, the rate of retrieval appears to be lower in FFNs, which suggests that other factors might come into play in building the hidden representation of words.
 However, in Fig.
-4(b) we plot the cumulative retrieval rate, i.e., for each layer, whether the word was identified in any layer so far.
-Our results indicate a different conclusion: the FFN update vectors match the (single-token) input representation of the word 70% of the time in layer, particularly both earlier and more frequently compared to the hidden representation.
-This suggest a larger variance in the exact layer in which FFNs identify the word compared to the hidden state.
-Our results indicate that the FFNs play a pivotal role in reconstructing word-level information from sub-word tokens.
+4(b) we plot the cumulative retrieval rate, i.e., for each layer, whether the word is identified in any layer so far.
+Our results indicate a different conclusion: the FFN update vectors match the (single-token) input representation of the word 70% of the time in any layer, particularly both earlier and more frequently compared to the hidden representation.
+This hints that FFNs indeed play a substantial role in building the internal word representations in LLMs.
+We further investigate the role FFNs play in detokenization through ablation experiments on FFN updates in Llama2-7B.
+Specifically, we test whether these updates are necessary for the word representation to emerge in the residual stream.
+To do so, we artificially split single-token words in a similar setting to Sec.
+4.1, this time focusing on derived words, formed by adding a suffix to a root word.
+[^10] We examine words with three common suffixes: “ing”, “ion”, and “est”.
+We split these back to two parts, for example, we divide “eating” to “eat” and “ing”, so that processing the suffix token is essential to reconstructing the word correctly.
+Using logit lens, we examine the FFN updates to the suffix token’s residual stream, and selectively ablate those associated with the original single-token word ($\sim$5of layers).
+As a control, for each word, we ablate an equal number of random FFN updates.
+Our results (Fig.
+12 in App.
+E) show that removing the updates carrying full-word representations dramatically reduces retrieval rates—from 85% without ablation to just 18%.
+In contrast, ablating random FFN updates has little to no effect.
+This indicates that FFN updates are essential for a detokenized representation to emerge in the final token.
+But does this affect the model’s ability to ”understand” the word, particularly when predicting the next token?
+To investigate this, we evaluate the model’s ability to retrieve the capital cities of countries, using the prompt “The capital of [COUNTRY] is ______”.
+We use single-token country names artificially split into two tokens, and follow the previous ablation protocol.
+We observe notable patterns: randomly ablating FFN updates reduces performance from 88% to 74%, [^11] We note that baseline performance when passing country names without artificial splits is 95%.
+likely due to disrupting the model’s factual recall mechanism (Meng et al., 2022b).
+However, specifically canceling updates detected as country’s name (similarly occurring in only $\sim$5of layers) causes a sharp drop to 41%.
+Overall, our results suggest that FFNs are central to detokenization, and actively reconstruct full-word meanings in LLMs.
 5.2 Token Aggregation
-Our results hint that LLMs retrieve words form the FFN layers, which allows the final token to represent these words.
-But this role of the FFN layers only starts to emerge in layers 3–4.
+Our results suggest that LLMs use the FFN layers to store and retrieve word representations, which are accessed in the final token to reconstruct the complete word.
+However, this role of the FFN layers only begins to emerge around layers 3–4.
 What happens earlier?
-We rely on previous work that showed that in the early layers the model focuses on integrating information from nearby tokens to compose entities (Lad et al., 2024), and hypothesize that the model starts by aggregating information from the previous sub-word tokens.
-(a) Word retrieval per layer
-(b) Cumulative word retrieval per layer
-Figure 4: Word retrieval from the FFN layers vs.
-from the hidden state.
-(4a) Retrieval rates across layers.
-The FFN values peak before word retrieval begins in the hidden layers.
-(4b) Its cumulative version, showing that word retrieval occurs earlier and more frequently in FFNs than in the hidden representation.
-To test this, we take all 2-token words in a subset of 1,000 documents from WikiText-103 (a total of 5,571 words), and feed them (along with their context) to the Llama2-7B model.
-We then measure the average attention weights of the final token to the prefix tokens in each layer.
-As control, we measure the average attention weights assigned by single-token words to their preceding token.
-Figure 5: Attention weights for 2-token words: Early peaks (layers 2–3) show high attention values from the 2nd sub-word token to the first, but these values decline rapidly.
+We build on previous work showing that early layers primarily integrate information from nearby tokens to compose entities (Lad et al., 2024), and hypothesize that the model starts by aggregating information from the previous sub-word tokens.
+Figure 5: Attention weights for 2-token words: Early peaks (layers 2–3) show high attention values from the second sub-word token to the first, but these values decline rapidly.
 Attention from single-token words to their previous token shows a similar trend, though with substantially lower values at first, which become higher later.
+To test this, we extract all two-token words in a subset of 1,000 WikiText-103 documents (a total of 5,571 words), and feed them, along with their context, to Llama2-7B.
+We then measure the average attention weights of the final token to the prefix token in each layer.
+As a control, we also measure the average attention weights assigned by single-token words to their preceding token.
 Our results (Fig.
-5) support previous findings (Lad et al., 2024)—the attention to previous tokens is high in the first 2–3 layers, but then declines sharply (up to $\sim$90decline).
-[^8] Experiments with 3- and 4-token words show a similar trend, see Appendix B.
-Considering the attention from single-word tokens to their previous token, we also observe this peak phenomenon, but importantly—here the initial peak is far lower than in our multi-token words [^9] The diverging patterns between single and multi-token words are statistically significant; see Appendix B..
-However, we also note that the following values are in turn higher.
-These results suggest that LLM strongly attend to their sub-word tokens at first, but then largely ignore them.
-Our results suggest that the model performs a detokenization process by first aggregating information from the prefix tokens into the final token’s hidden representation.
-It then refines the representation of the final token in the FFN layers to retrieve the full word.
-This two-stage process of token aggregation and concept retrieval provides insight into the internal mechanisms of LLMs and how they handle sub-word tokens to reconstruct word-level representations.
+5) support previous findings (Lad et al., 2024)—the attention to previous tokens is high in the first 2–3 layers, but then declines sharply (by up to $\sim$90%).
+[^12] Experiments with 3- and 4-token words show a similar trend, see App.
+B.
+For single-token words, we observe a similar attention pattern to the previous token in the first layers, but importantly—the initial peak is significantly lower than in multi-token words.
+[^13] The diverging patterns between single and multi-token words are statistically significant; see App.
+B.
+Still, in later layers, the attention weights of single-token words are in turn higher than for multi-token words.
+These results suggest LLMs strongly attend to the preceding sub-word tokens of multi-token words at first, but then largely ignore them.
+Altogether, our results suggest that LLMs perform a detokenization process by first aggregating information from the prefix tokens into the final token’s hidden representation, and then refining the representation of the final token using the FFN layers to retrieve the full word’s concept representation.
+This two-stage process of token aggregation and concept retrieval provides insight into the mechanisms LLMs use to handle sub-word tokens and reconstruct word-level representations.
 6 Expanding LLM Vocabulary Without Finetuning
-We have shown that language models internally fuse multi-token words into single-token representations, which models can decode to generate the complete word.
-This raises the question whether models can make more general use of such fused representations when processing textual input, by using them in place of the original, multi-token embeddings.
-Conversely, models were shown to foreshadow several future tokens in a single hidden state without being explicitly trained to do so (Pal et al., 2023); can we leverage the fused representations to enable models to generate multi-token words in a single inference step?
-We next explore whether such vocabulary expansion is possible without any updates to the model’s parameters.
-[^10] We do add new entries to the model’s embedding and unembedding matrices, but these are inherently required to expand the vocabulary and represent the new vocabulary words.
-The potential of such methods is high, as even tokenizers intentionally trained for multilingual support result in extremely longer tokenization lengths for non-English languages (up to 13x for some languages), affecting cost and latency (Ahia et al., 2023; Sengupta et al., 2023; Petrov et al., 2024).
-Still, previous attempts to expand tokenizer vocabulary are limited in number, and require further training on large corpora (Kim et al., 2024; Zhao et al., 2024) or attempt to force alignment between languages in embedding space (Wang et al., 2019), rather than rely on the internal detokenization mechanism of LLMs.
-We first experiment with expanding the input embeddings matrix $E$, leaving the output unembedding matrix $U$ unchanged.
-To extract fused representations, we begin with the same approach as in Section 4.2: Given a multi-token word, we run it through the model using a prompt template $P$ [^11] We use the template “x x x x” where x is the new vocabulary word; preliminary experiments showed it provides better results than the template “x”, possibly due to the effect a token’s position has on its representation.
-We then identify the first layer in which the last token is successfully decoded into the full word (Ghandeharioun et al., 2024).
-We extract the hidden state of the last token in that layer, and add it as the word’s fused representation to $E$.
-If the word is never successfully decoded from the representation in any of the layers, we don’t attempt to add it to the vocabulary, as we assume it is not in the model’s inner lexicon.
-We experiment with two approaches: identity, where we use the fused representation as is; and linear, where we first transform it with a linear map $T_{\ell,E}$.
-We learn $T_{\ell,E}$ by fitting a linear regression from the layer $\ell$ hidden states of all single-token words when passed with the template $P$, to their corresponding input embeddings.
-Importantly, $T_{\ell,E}$ is both not part of the original model, and does not depend on the new, originally multi-token words we add to the vocabulary.
-Original vocabulary tokens Newly added words
-Top-1 Acc. Mean Reciprocal Rank Top-1 Acc. Mean Reciprocal Rank
-Original vocabulary 57.7% 0.679 47.5% 0.556
-Last token embedding ($E$) 54.5% 0.647
-Fused-Identity ($E$) 53.5% 0.638
-Fused-Linear ($E$) 56.7% 0.669
-Fused-Linear ($E$ + $U$) 56.7% 0.667 19.5% 0.598
-Table 1: Performance of finetuning-free vocabulary expansion for Llama2-7B on WikiText-103.
-Different rows correspond to baselines (first two rows) and proposed approaches (last three).
-Methods with $E$ add new multi-token words as single-token entries in the embedding matrix, while those with $U$ add new entries to the unembedding matrix.
-Column groups correspond to evaluations of language model performance on different token groups: those already in the vocabulary (left) and the newly added entries (right, relevant only when expanding $U$).
-Results marked with $\ast$ are for predicting the 1st token of a multi-token word (rather than predicting the new entry of the complete word).
-Our best approach (Fused-Linear) leverages the fused representations emerging within models after detokenization, to add new vocabulary entries for multi-token words.
-We show it maintains performance on the original vocabulary, and enables the model to successfully utilize entries in $U$ that were never observed during training.
-We apply our approach to expand the vocabulary of Llama2-7B with all multi-token words found in the WikiText-103 test set, resulting in 4.3k newly added words.
-To examine whether the model can correctly use the new entries, we evaluate it in a language modeling setup, when all newly added words are replaced with their fused single-token embeddings, similar to our setup in Section 4.2 (see righthand side in Fig.
-1).
-We compare our approach to two baselines: using the original vocabulary without any new entries; and using the unprocessed last token embedding of each new word as its new vocabulary entry (instead of the fused representations).
-For evaluation, we cannot use perplexity, the standard LM measure, as the different approaches have different vocabularies, which influences perplexity scores and prevents a fair comparison.
-Instead, we measure the token-level top-1 accuracy and the mean reciprocal rank of the ground truth tokens; naturally, we do not report the ranking of sub-word tokens forming the newly added words, as they only serve as ground truth in our baseline.
-We report our results in Table 1.
-Examining model performance when using the fused representations in the identity approach, we find that while the model can successfully decode the full word information when explicitly prompted to do so, it fails to utilize the fused representation for inference, and achieves slightly worse performance than when using the unprocessed last token embeddings.
-However, using the linear approach by first applying $T_{\ell,E}$ transforms the fused representation into a usable form for initial layers, as we find the model can then utilize the newly added words and maintains its performance.
-[^12] A possible explanation for the importance of $T_{\ell,E}$ is that input embeddings have very different distributions from intermediate hidden states, as LLMs are extremely sensitive to 1st layer ablations (Lad et al., 2024).
-Given the model’s ability to use fused representations as inputs, we next enable the model to generate multi-token words in a single inference step, by adding new entries to the unembedding matrix $U$ derived from their fused representations.
-In light of our previous results, we only experiment with the linear approach.
-We follow the same steps as before, this time applying linear maps $T_{\ell,U}$ which we learn by fitting linear regressions from the layer $\ell$ hidden states of all existing single-token words, to their corresponding rows in the unembedding matrix.
-We evaluate whether the model chooses to generate the new vocabulary entries, or rather the first token of the original multi-token words (as it does when using the unedited vocabulary).
-Surprisingly, we find the model assigns, on average, higher rank to the new vocabulary entries over their original 1st sub-word tokens in the appropriate contexts (Table 1, last row).
-While the top-1 accuracy is higher for the 1st sub-word tokens in the unedited vocabulary, we find that when new vocabulary entries are not assigned high ranks, the original sub-word will often be predicted.
-We further observe that in 7.1% of examples, the new entry achieves the top-1 rank even though the original sub-word in the baseline does not.
-Notably, this also does not compromise performance on existing vocabulary words.
-Our results suggest that the model can sometimes try to generate the fused representation, but as it does not exist in its (original) unembedding vocabulary, the first sub-word token is selected instead.
-Overall, our findings show potential to substantially decrease inference times for LLMs.
-In future work, we will explore their application to multilingual LLMs (Petrov et al., 2024).
+We have shown that language models internally fuse multi-token words into a single-token representation, and that they can further “read” these representations as inputs, and decode the original multi-token words.
+This raises the question: can models use these fused representations instead of the original multi-token inputs—to encode input prompts using less tokens and reduce computation?
+Similarly, models were shown to implicitly predict several future tokens in a single hidden state without being explicitly trained to do so (Pal et al., 2023); can we leverage these representations to enable models to predict multi-token words in a single inference step?
+Motivated by our findings, we explore whether we can expand the model’s vocabulary with new input and output embeddings for originally multi-token words, without any updates to model parameters.
+[^14] We do add new entries to the model’s embedding and unembedding matrices, but these are intrinsically required to expand the vocabulary and represent the new vocabulary words.
+This goal is of practical importance, especially for low-resource languages and domains: even tokenizers built for multilingual support often produce substantially longer token sequences for non-English languages—up to 13 times longer than equivalent English texts—impacting inference cost and speed (Ahia et al., 2023; Sengupta et al., 2023; Petrov et al., 2023).
+Still, prior attempts to expand tokenizer vocabulary post-hoc are limited in number, and require substantial additional training (Kim et al., 2024; Zhao et al., 2024).
+In contrast, we propose to detect words the model “knows” and successfully detokenizes to a single vector, and use this to obtain new token embeddings.
+Figure 6: Our 3-step method to expand LLM vocabulary without updates to core model parameters.
+Our framework for vocabulary expansion follows a 3-step process (Fig.
+6).
+Given a multi-token word $w$ that we would like to add to the vocabulary, and the model’s original input embedding and output unembedding matrices $E$ and $U$ with hidden dimension $d$, we (1) extract a detokenized, single-token representation $r$ for $w$ using a similar approach to Sec.
+4.2: pass $w_{i}$ as input to the model and apply Patchscopes (Ghandeharioun et al., 2024) with prompt $P$ [^15] For $P$, we use the template “x x x x” where x is the patched representation of the new word.
+to the last token’s hidden states at all layers.
+We then identify $\ell$, the earliest layer at which the hidden state is successfully decoded into the full word, and set $r$ as that hidden state.
+Next, we (2.1) learn a set of linear maps $T_{\ell,E}$ and $T_{\ell,U}$ to project hidden states from layer $\ell$ of the model to the embedding and unembedding spaces, based only on the existing in-vocabulary tokens.
+[^16] We learn $T_{\ell,E}$ and $T_{\ell,U}$ by fitting an orthogonal procrustes transformation from the layer $\ell$ hidden states of all in-vocabulary tokens (when passed as inputs on their own), to their corresponding embeddings or unembeddings (see App.
+G for details).
+Importantly, these projections only rely on the model’s existing vocabulary, and do not depend on which multi-token words we choose to add to the vocabulary.
+Then, (2.2) for each detokenized representation $r$ taken from layer $\ell$, we apply $T_{\ell,E}$ and $T_{\ell,U}$ to obtain the initial entries in the embedding and unembedding matrices to represent the word, $\hat{e}$ and $\hat{u}$.
+Finally, after computing the initial entries for all new words, we (3) refine the new representations to obtain the final entries: we initialize two $d\times d$ all-zeros matrices $W_{E}$ and $W_{U}$, and set the new entries as $e=\hat{e}+W_{E}\hat{e}$ and $u=\hat{u}+W_{U}\hat{u}$.
+We train the refinement matrices $W_{E}$ and $W_{U}$ jointly in a short continued pretraining run, while keeping all other parameters frozen.
+Finally, we compute the final $e$ and $u$, and use these to represent the new word in the expanded vocabulary.
+Importantly, if a word is never successfully decoded from the hidden states in any of the layers in step (1), we assume it is not found in the model’s inner lexicon, and therefore do not add it to the vocabulary.
+We apply our approach to Llama2-7B and experiment with three datasets: WikiText-103 (Merity et al., 2017), abstracts of biomedical articles from PubMed (Xiong et al., 2024), and the Arabic split of Wiki40B (Guo et al., 2020).
+For each dataset, we expand the model’s vocabulary with all multi-token words that appear at least $m$ times in the test set.
+[^17] We use $m=1,5,50$ for WikiText-103, PubMed and Arabic Wiki40B respectively.
+We then learn the refinement matrices $W_{E}$ and $W_{U}$ using 20M tokens from the train set.
+[^18] We use a sequence length of 512 and train on 10,000 sequences, taking up to 30 minutes on a single GPU.
+We evaluate models in a next-word prediction setup and compare our approach against two baselines: the original model with an unmodified vocabulary (original); and an expanded model that follows our framework but uses the mean embeddings of the word’s tokens in $E$ and $U$ to initialize its new representation, instead of the detokenized representations $r$, following Gee et al.
+(2022; mean embedding).
+For evaluation, perplexity is not a suitable metric, as differences in vocabulary between the approaches skew perplexity scores and prevent a fair comparison.
+Instead, we measure the model’s token-level top-1 accuracy when it is given each token’s previous context.
+WikiText-103 PubMed Wiki40B-Arabic
+Original Mean Embed. Ours Original Mean Embed. Ours Original Mean Embed. Ours
+New words new token 0.071 0.171 0.123 0.180 0.117 0.402
+original or new token 0.322 0.178 0.284 0.280 0.171 0.259 0.413 0.119 0.414
+All words 0.522 0.473 0.519 0.517 0.479 0.511 0.535 0.211 0.532
+Table 1: Token-level accuracy of finetuning-free vocabulary expansion for Llama2-7B on three datasets.
+We compare our method using detokenized representations (Ours) to the original model (Original) and to an expansion baseline using the word’s average token embeddings from $E$ and $U$ (Mean Embed.; Gee et al.
+2022).
+Accuracy is reported for newly added words, where we distinguish between correctly predicting the new tokens (top row) and predicting either the word’s original first token, as the unexpanded model would, or its new token (middle row).
+We also report overall performance on all tokens (last row).
+Our method enables the frozen model to use the newly added input and output embeddings, while maintaining overall model performance.
+Our results (Table 1) indicate that models can generalize to new vocabulary entries surprisingly well when these are initialized with their own detokenized representations.
+Unlike the mean embeddings expansion baseline, which both struggles with using new tokens and degrades overall performance, our method allows the model to successfully integrate new vocabulary words while preserving its accuracy on existing words.
+This effect is particularly pronounced in Arabic Wiki40B, where the model almost always selects new tokens instead of the original tokenization (40.2% vs 41.3%), demonstrating the potential of our approach for multilingual and domain-specific applications.
+Our method also shows potential to reduce the number of tokens processed during encoding and inference: across the three datasets, we observe a reduction in the total number of tokens processed during encoding by 10.5% to 14.5% (see Table 5 in App.
+H), while maintaining model performance.
+In future work, we will further explore the application of our framework to multilingual adaptation (Petrov et al., 2023; Alabi et al., 2022) and continual domain-adaptive pretraining (Ke et al., 2023; Yıldız et al., 2024; Gururangan et al., 2020).
 7 Conclusion
-The ability of LLMs to comprehend and generate language relies on intricate internal processes, and understanding these mechanisms is crucial for improving model performance as well as making them more efficient.
+The ability of LLMs to comprehend and generate language relies on intricate internal processes, and understanding these mechanisms is crucial for improving model performance and efficiency.
 In this work, we unraveled the word detokenization process, shedding light on how models internally transform fragmented sub-word tokens into coherent word representations formed at the last token.
 Our results indicate that this mechanism manifests in early to middle layers, where models attempt to reconstruct words by mapping them to an inner lexicon using their FFN layers.
-We provided evidence this lexicon is more exhaustive than the tokenizer’s vocabulary, and could facilitate the model’s ability to recognize words even amidst noise.
+We provided evidence this lexicon is more exhaustive than the tokenizer’s vocabulary, and could help models to recognize words even amidst noise.
 Our work also unlocks practical avenues for optimizing tokenization, as well as the speed and cost of inference.
@@
 We hope our work paves the way for more efficient and versatile models.
-Acknowledgements
-This work was supported in part by the Israel Science Foundation (grant no. 2045/21) and by NSF-BSF grant 2020793. In addition, we thank Iddo Yosha for his valuable feedback.
+Acknowledgments
+This work was supported in part by the Israel Science Foundation (grant no.
+2045/21) and by NSF-BSF grant 2020793.
+We are grateful to Iddo Yosha for his valuable insights and thank the anonymous reviewers for their constructive feedback.
 Appendix A Words vs. Nonwords Analysis
-To further analyze results from Section 3, we conduct a failure analysis, focusing on false negatives (FN) and false positives (FP).
-The former, valid words misclassified as gibberish, are often rare or complex words, suggesting that the model’s internal vocabulary may lack representations for less common words.
-On the other hand, false positives—where gibberish is misclassified as a word—typically involved nonwords that closely resemble valid words, likely due to shared sub-word structures.
-See Table 2 for a few examples.
-(a) accuracy postfix and last token
-(b) positive predictions over prefix and last token
-Figure 6: Comparisons between last token and the prefixes.
-(Left) we see that the accuracy over the prefix tokens reaching only 61% vs the 89% with the last token’s representation.
-(Right) a more detailed statistics over the positive predictions.
-The TP rate over the last token’s representations reaching 84% while reaching only 43% for the prefix ones
+To further analyze our results from Sec.
+3, we conduct a failure analysis, focusing on false negatives (FN) and false positives (FP).
+Our analysis indicates that the former—valid words misclassified as gibberish—are often rare and complex words, suggesting that the model’s internal vocabulary may lack representations for infrequent words.
+On the other hand, false positives—where gibberish is misclassified as a word—typically involves nonwords that closely resemble valid words, likely due to shared sub-word structures.
+See Tab.
+2 for a few examples.
 Original Word Predicted Value True Value Status
@@
 _nonwith word gibberish FP
-Table 2: Examples of false negative (FN) and false positive (FP) for the word vs. nonword experiment (Section 3).
+Table 2: Examples of false negative (FN) and false positive (FP) for the word vs. nonword experiment (Sec. 3).
+We further conduct additional experiments to compare the performance of our custom nonword dataset with the linguistically-motivated ARC Nonword Database (Rastle et al., 2002), which contains morphologically plausible nonwords designed to resemble real words to humans.
+Our results (Fig.
+7) indicate a higher differentiability on the ARC dataset compared to our own dataset, which highlights that our nonword creation procedure effectively mitigates potential biases.
+This result strengthens our conclusions: the last token provides significant cues for distinguishing words from nonwords, particularly in the middle layers of the model after several layers of processing.
+Figure 7: Accuracy comparison by layer for distinguishing words from nonwords.
+The ARC Nonword Database (blue line) exhibits the highest accuracy across layers, potentially due to its morphologically plausible nonwords.
+In contrast, our dataset (orange line, last token) achieves slightly lower accuracy, demonstrating the effectiveness of our bias-mitigated nonword generation process.
+The penultimate token from our dataset (green line) shows significantly lower accuracy, highlighting the importance of the last token in distinguishing words from nonwords.
+In another experiment, we evaluate the penultimate token representation against nonword tokens for words of length three or more tokens.
+For example, in a word like “unhappiness,” we use the inner token “h” (e.g., un-h-appiness).
+Unlike the final tokens, penultimate tokens are poorly distinguishable from nonwords, achieving significantly lower accuracies.
+These results suggest that the model’s ability to separate words from nonwords is highly dependent on the position and completeness of token representations.
+Finally, we expand the comparison of our dataset across three additional models—Llama3-8B, Mistral-7B, and Yi-6B.
+Our results across these models (Fig.
+8) mirror the trends observed with Llama2-7B, with accuracy improving as layers deepened but peaking and gradually declining in later layers.
+This consistency further supports our findings that the model’s ability to classify words vs.
+nonwords relies on nuanced patterns in internal token representations, unaffected by co-occurrence biases or morphological artifacts.
+Figure 8: Mean true status accuracy by layer for distinguishing words from nonwords across multiple models (Llama2-7B, Llama3-8B, Mistral-7B, Yi-6B) using our dataset.
+All models exhibit similar trends: accuracy improves across initial layers, peaks in the middle layers, and declines in deeper layers.
+This pattern demonstrates the robustness of our findings across different model architectures and sizes.
 Appendix B Statistical Analysis of Token Aggregation in Multi-token Words
-We repeat the multi-token experiments Section 5.2 for 3- and 4-token words. Results Fig. 7 show a very similar trend to the 2-token words (Fig. 5).
-We present the detailed statistical analysis conducted of our experiments in Section 5.2 to examine token aggregation in multi-token words compared to single-token words (control group).
+We repeat the multi-token experiments Sec. 5.2 for 3- and 4-token words. Our results (Fig. 9) show a very similar trend to 2-token words (Fig. 5).
+We further present the detailed statistical analysis conducted of our experiments in Sec.
+5.2 to examine token aggregation in multi-token words compared to single-token words (control group).
 The objective is to investigate whether the model exhibits significantly different attention patterns between these groups, indicative of the detokenization process.
@@
 Table 3: Results of one-sided t-tests comparing attention weights between multi-token and single-token words across layers over Llama2-7B model.
-To determine significance, we performed two one-sided t-tests per layer for the Llama2-7B case: (1) Testing whether attention to prefix tokens in multi-token words is higher than to previous tokens in single-token words.
+To determine significance, we perform two one-sided t-tests per layer for the Llama2-7B case: (1) Testing whether attention to prefix tokens in multi-token words is higher than to previous tokens in single-token words.
 (2) Testing whether attention to previous tokens in single-token words is higher than to prefix tokens in multi-token words.
-Table 3 shows the p-values and significance levels for each layer. Significance levels are denoted as ns (not significant), and *** ($p<0.001$).
+Table 3 shows the p-values and significance levels for each layer.
+Significance levels are denoted as ns (not significant), and *** ($p<0.001$).
 (a) Llama2-7B Attention weights for 3-tokens words
@@
 From Layers 3 to 17, attention to single-token words is higher, indicating a shift in attention focus from prefix attention in relation regular close tokens.
-Notably, there are intermittent increases in attention to prefix tokens at Layers 18, 21, 25, 27, 29, and 30, possibly signaling a transfer into a prediction ensembling phase in which the attention in general is less important (Artzy & Schwartz, 2024) and therefore we don’t see a coherent pattern.
-In conclusion, the attention mechanisms differ significantly between multi-token and single-token words, confirming that the detokenization process involves initial amplification of attention to sub-word tokens followed by a reduction as the model processes word representations.
-Appendix C Word Retrieval for Single-Token and Multi-Token Words Across All Models Comparison
-This section presents a detailed comparison of word retrieval performance (Section 4) across several models (Llama2-7B, Llama3-8B, Yi-6B, and Mistral-7B) for both single-token and multi-token words.
+Notably, there are intermittent increases in attention to prefix tokens at Layers 18, 21, 25, 27, 29, and 30, possibly signaling a transfer into a prediction ensembling phase in which the attention in general is less important (Ben Artzy & Schwartz, 2024) and therefore we don’t see a coherent pattern.
+In conclusion, the attention mechanism differs significantly between multi-token and single-token words, indicating that the detokenization process involves initial amplification of attention to sub-word tokens followed by a reduction as the model obtains whole-word representations.
+Appendix C Word Retrieval for Single-Token and Multi-Token Words Across Models
+This section presents a detailed comparison of word retrieval performance (Sec.
+4) across several models (Llama2-7B, Llama3-8B, Yi-6B, and Mistral-7B) for both single-token and multi-token words.
 The evaluation focuses on how effectively each model retrieves the original word across layers, especially in challenging cases like artificially separated single-token words, typos, and multi-token words.
+Across all experiments (Fig.
+10), we observe a similar trend where the retrieval rate increases over the first several layers, peaks around the middle layers, and then decreases in the later layers.
+The main difference across models lies in the peak performance, especially in cases involving typos and multi-token words, where more advanced models such as Llama3-8B and Mistral-7B demonstrate superior performance in reconstructing the original word representations.
 Figure 10: Layer-wise word retrieval rates for single-token and multi-token words across all models.
-Across all experiments, we observe a similar trend where the retrieval performance increases over the first several layers, peaks around the middle layers, and then decreases in the later layers.
-The main difference across models lies in the peak performance, especially in cases involving typos and multi-token words, where more advanced models such as Llama3-8B and Mistral-7B demonstrate superior performance in reconstructing the original word representations.
-Appendix D Cumulative Word Retrieval for Single-Token and Multi-Token Words
-In this section, we present the results of cumulative word retrieval (Section 4.1) across various models, focusing on both single-token and multi-token words.
+C.1 Cumulative Word Retrieval for Single-Token and Multi-Token Words
+Fig.
+11 shows the results of cumulative word retrieval (Sec.
+4) across various models, focusing on both single-token and multi-token words.
 For each model, we analyze the ability of the LLM to retrieve the original word from sub-word tokens across its layers.
 The cumulative retrieval is calculated as the proportion of words that are successfully retrieved at each layer, with the percentage increasing as more words are recovered throughout the model’s layers.
-Single-Token Words: The cumulative word retrieval for single-token words—those that are artificially split into sub-word tokens (via typographical errors or manual splits)—shows a rapid increase in retrieval success in the early layers.
+Single-token words
+The cumulative word retrieval for single-token words—those that are artificially split into sub-word tokens (via typographical errors or manual splits)—shows a rapid increase in retrieval success in the early layers.
 For Llama2-7B, for instance, cumulative retrieval reaches 93.2% for words split by manual intervention, and 66% for words affected by typos by the middle layers.
 This pattern is observed across models, with retrieval generally peaking around layers 15-20.
-Multi-Token Words: For multi-token words, which are naturally split due to being out-of-vocabulary for the tokenizer, the cumulative retrieval process follows a similar trajectory.
+Multi-token words
+For multi-token words, which are naturally split due to being out-of-vocabulary for the tokenizer, the cumulative retrieval process follows a similar trajectory.
 However, in models like Llama2-7B, the retrieval peaks earlier in the model, with a cumulative retrieval rate of 77.41%.
@@
 Figure 11: Cumulative word retrieval for single-token and multi-token words across all models.
-Appendix E Introducing Typos for Single-Token Words
-In this section, we describe the process of introducing typos into single-token words to split them into multiple tokens (Section 4.1).
+Appendix D Introducing Typos for Single-Token Words
+In this section, we describe the process of introducing typos into single-token words to split them into multiple tokens (Sec.
+4.1).
 The modification applies to words longer than four characters and involves randomly performing one of three operations: substituting two characters, deleting a character, or inserting a new character.
@@
 Table 4 shows examples of the different splits.
-Description perturbed . new tokens
+Description perturbed new tokens
 Substitution of two characters develpoment [‘de’, ‘vel’, ‘p’, ‘oment’]
@@
 Insertion of one character develfopment [‘dev’, ‘elf’, ‘op’, ‘ment’]
-Table 4: Examples of the different typos we consider, exemplified by perturbing the single-token word “development” (Section 4.1).
+Table 4: Examples of the different typos we consider, exemplified by perturbing the single-token word “development” (Sec. 4.1).
+Appendix E Ablation Experiment on Suffix-split Words
+In Sec.
+5, to test the role of FFN layers in detokenization, we conduct an intervention-based experiment measuring how ablations to FFN updates affect word retrieval.
+We run this experiment on all single-token words from WikiText-103 that end with one of three common suffixes: “ing,” “ion,” or “est”.
+Each word is then artificially split to two parts—the root word and the suffix.
+For example, the word “eating” is split into “eat” and “ing”, while “connection” is split into “connect” and “ion”.
+This ensures that processing the suffix token is necessary to reconstruct the identity of the word, and reduces possible effects of strong distributional artifacts of token co-occurrence.
+Using logit lens, we identify FFN layers where the update to the residual stream can decoded as the original single-token word.
+We then ablate these layers by zeroing out their updates to the residual stream.
+As a control, we ablate the same proportion (5%) of random FFN layers.
+Our results, shown in Fig.
+12, reveal that ablating the identified FFN layers lead to a sharp drop in word retrieval rates, effectively disrupting the detokenization process.
+In contrast, ablating random layers has little to no effect on retrieval accuracy.
+This indicates that the FFN layers play a critical role in reconstructing word representations rather than simply enhancing contextualization.
+Figure 12: Comparison of retrieval rates using logit lens for suffix-split words under three conditions: ablation of identified FFN layers (blue line), regular retrieval (orange line), and control with random FFN layer ablation (green line).
+Ablation of critical layers causes a sharp drop in retrieval accuracy, highlighting their importance in detokenization.
+Appendix F Comparision Between Logit Lens and Cosine Similarity
+To validate our use of the logit lens, we repeat the artificial split-word experiment using cosine similarity. We use the same setting of artifical splits based on suffixes detailed above in App. E.
+Figure 13: Comparison of retrieval accuracy for split-word experiments using either logit lens (blue line) and cosine similarity (orange line) across model layers.
+For the logit lens, we adapt the standard implementation to measure similarity between the hidden representation of the final token and the input embedding space (vocabulary space), rather than the output embedding space typically used to predict the next token.
+This adjustment allows us to directly evaluate the alignment between the hidden states and the embeddings of the original words in the vocabulary.
+In parallel, cosine similarity is used to compute the similarity between the same hidden states and the embeddings of the original words in the vocabulary space.
+Our results, shown in Fig.
+13, demonstrate that both methods produce nearly identical patterns.
+In both cases, retrieval accuracy peaks in the middle layers, where the hidden representation of the suffix token aligns most closely with the original word.
+Beyond these middle layers, retrieval accuracy declines, likely reflecting the model’s shift toward representing the prediction of the next token rather than maintaining the full representation of the current word.
+These findings reinforce the validity of using the logit lens when adapted to the input embedding space, as a streamlined approach for analyzing model representations.
+This approach results in similar trends to cosine similarity while offering a more interpretable and direct framework for measuring alignment with the vocabulary space.
+Appendix G Learning the Linear Maps $T_{\ell,E}$ and $T_{\ell,U}$
+To expand the model’s vocabulary without modifying its core parameters, we construct linear transformations that map hidden states at different layers to the model’s embedding and unembedding spaces.
+By learning these transformations solely from the model’s existing vocabulary, we can infer new token representations without modifying any of the model’s core weights.
+This section details our method for learning these mappings.
+G.1 Extracting Model Representations
+Given a pretrained language model with input embedding matrix $E\in\mathbb{R}^{V\times d}$ and output unembedding (LM head) matrix $U\in\mathbb{R}^{V\times d}$, where $V$ is the vocabulary size and $d$ is the hidden dimension, we extract the representations used for learning the transformations as follows:
+  • Embedding and Unembedding Matrices: We extract the model’s embedding matrix $E$ and LM head matrix $U$ before making any modifications to the vocabulary.
+  • Hidden States Across Layers: For each token $t$ in the model’s original vocabulary, we pass $t$ as a single-token input and record its hidden state at every layer of the model.
+Let $h_{\ell}(t)$ denote the hidden state of token $t$ at layer $\ell$.
+G.2 Learning the Linear Mappings
+For each layer $\ell$, we aim to learn two linear transformations:
+  • $T_{\ell,E}$ that maps hidden states to input embeddings.
+  • $T_{\ell,U}$ that maps hidden states to unembedding representations.
+We learn these mappings as orthogonal Procrustes problems (Schönemann, 1966), which seek to find the best orthogonal transformation aligning two sets of vectors.
+Specifically, for each layer $\ell$, we solve:
+[…]
+[…]
+where $e_{t}$ and $u_{t}$ are the embedding and unembedding vectors of token $t$, respectively.
+The constraints enforce that each transformation preserves distances and does not distort the structure of the space.
+In our experiments, we use the Python implementation of Meng et al.
+(2022a).
+G.3 Normalization with RMS Scaling
+To preserve the relative magnitudes of embedding and unembedding entries, we normalize all representations using their root mean square (RMS) norm (Zhang & Sennrich, 2019). Specifically:
+  1. 1.
+    Preprocessing Training Representations: Before fitting the Procrustes transformations, we normalize all hidden states, embeddings, and unembedding vectors by dividing each vector $x$ by its RMS norm:
+    $x_{\text{norm}}=\frac{x}{\|x\|_{\text{RMS}}},\quad\text{where }\|x\|_{\text{RMS}}=\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}}.$ (3)
+  2. 2.
+    Applying the Learned Maps: When using the trained transformations on new detokenized representations $r$, we apply the following steps:
+    1. (a)
+      Normalize $r$ by its RMS norm: $r_{\text{norm}}=\frac{r}{\|r\|_{\text{RMS}}}$.
+    2. (b)
+      Apply the learned transformation: $\hat{e}=T_{\ell,E}r_{\text{norm}}$, $\hat{u}=T_{\ell,U}r_{\text{norm}}$.
+    3. (c)
+      Rescale by the mean RMS of the target space:
+      $e=\hat{e}\cdot\mathbb{E}[\|e_{t}\|_{\text{RMS}}],\quad u=\hat{u}\cdot\mathbb{E}[\|u_{t}\|_{\text{RMS}}].$ (4)
+This ensures that the new embeddings and unembeddings maintain a scale consistent with the original vocabulary.
+Appendix H Efficiency Gains from Vocabulary Expansion
+Beyond improving model performance on newly added words, our vocabulary expansion method directly reduces the number of tokens required to encode input text, and could lead to further potential efficiency gains in inference.
+Dataset # Attempted # New Words Token Reduction
+WikiText-103 14.1k 10.1k 10.5%
+PubMed 9.5k 5.4k 13.5%
+Wiki40B-Arabic 4.4k 0.7k 14.5%
+Table 5: Reduction in average sequence length when encoding text with the expanded vocabulary for Llama2-7B.
+# Attempted is the number of multi-token words tested for expansion, while # New Words is those detected as detokenized using Patchscopes.
+Table 5 summarizes the average reductions in sequence length achieved by encoding texts from each domain using the expanded vocabulary instead of the original vocabulary in the three datasets.
+Using our method, we attempt to expand Llama2-7B’s vocabulary with all multi-token words appearing at least $m$ times in the test set, where we use $m=1,5,50$ for WikiText-103, PubMed, and Arabic Wiki40B, respectively.
+We find that the token reduction rates depend on the success rate of extracting detokenized representations using Patchscopes (which in turn depend on the textual domain and language).
+WikiText-103 achieves a 10.5% reduction in token count, with 72.9% of attempted words successfully converted into new vocabulary entries.
+PubMed shows a higher token savings rate (13.5%) despite a lower success rate (58.3%), as its vocabulary expansion targets more domain-specific multi-token terms.
+The largest efficiency gain is observed in Arabic Wiki40B, where encoding with the expanded vocabulary reduces token count by 14.5%, highlighting the method’s potential for languages with inherently longer token sequences.
+However, the success rate of identifying detokenized representations in Wiki40B is lower (16.0%), suggesting room for improvement in expanding vocabularies for morphologically rich languages, and suggesting many Arabic words are not stored in the model’s inner lexicon.
+These results suggest that post-hoc vocabulary expansion can significantly reduce the computational cost of inference, particularly for non-English languages and domain-specific texts, without requiring any modifications to the model’s core parameters.